### v1

# Addressing Model and Data Heterogeneity in Multimodal Large Language Model Training

Abstract
Multimodal large language models (LLMs) have demonstrated significant potential in a wide range of AI applications. Yet, training multimodal LLMs suffers from low efficiency and scalability, due to the inherent model heterogeneity and data heterogeneity across different modalities.
We present MMScale, an efficient and adaptive framework to reform the training of multimodal large language models on large-scale clusters. MMScale exploits the system characteristics of multimodal LLM training to achieve high efficiency and scalability. The core of MMScale is the adaptive resource allocation and data-aware reordering techniques to eliminate the model and data heterogeneity respectively. We also tailor system optimizations for multimodal LLM training to offload certain operations from the GPU training. We evaluate MMScale across different sizes of multimodal LLMs on a large-scale production cluster with thousands of GPUs. The experimental results show that MMScale achieves 54.7% Model FLOPs Utilization (MFU) when training a 72B multimodal LLM on 1172 GPUs and outperforms Megatron-LM by up to 2.2 on throughput. The ablation study shows the main techniques of MMScale are both effective and lightweight.
1 Introduction
Recent advances in large language models (LLMs) are catalyzing a new wave of AI applications [1]. However, LLMs are predominantly text-based, restricting their ability to understand and generate multimodal content. Emerging multimodal LLMs address this gap by integrating various modalities such as texts, images, and audios into LLMs which significantly enhances LLM’s applicability. Multimodal LLMs demonstrate great potential in tasks like image understanding [2, 3, 4, 5], audio comprehension [6, 7], and embodied AI [8, 9]. Many organizations are actively developing their multimodal LLMs, such as OpenAI’s GPT-4o [10], Google’s Gemini [11] and PaLM-E [9], Meta’s Chameleon [12], etc.
Training multimodal LLMs demands vast computational resources. According to the scaling law [13], model size and training data volume are crucial for determining model capabilities. Substantial efforts are invested in training models with billions of parameters on trillion-scale tokens. For instance, Meta’s Chameleon [12] is a 34B multimodal LLM trained on more than 4.8 trillion tokens. Leading organizations often deploy large-scale clusters, equipped with thousands of GPUs, for such training tasks. One technical report [14] reveals that training a GPT-3 175B with 300 billion tokens costs $4.6 million and lasts one year using 355 V100 GPUs. Consequently, it is crucial to develop an efficient and scalable training framework to minimize costs and accelerate training.
Figure 1 depicts the mainstream model architecture of multimodal LLMs, comprising three primary modules: the modality encoder, LLM backbone, and the modality generator [15, 16]. These modules are linked by the projector, which may incorporate MLP or cross-attention layers. The modality encoder transforms input from various modalities into a unified embedding space. The embedding vectors are then analyzed by the LLM backbone, a transformer model, to discern data patterns and inter-modal relationships. Subsequently, the modality generator translates this processed information back into coherent outputs tailored to each modality.
Megatron-LM [17] is the state-of-the-art framework for training LLMs, which comprise a stack of homogeneous transformer layers. It can be extended to train multimodal LLMs by treating the multimodal modules as additional layers within the LLM. However, training multimodal LLMs with Megatron-LM poses two substantial challenges: model heterogeneity and data heterogeneity. The fundamental issue of model heterogeneity stems from the need to process diverse modalities with different modules that vary dramatically in size and operator complexity. The model heterogeneity across different modules (i.e., modality encoder, LLM backbone, and modality generator) introduces severe pipeline bubbles, resulting in poor GPU utilization. Meanwhile, data heterogeneity bursts onto the scene due to the intricate and unstructured nature of multimodal input data. The data heterogeneity across different modality data leads to inter-microbatch and intra-microbatch training stragglers, which prolong the training duration and exacerbate the pipeline bubbles. These challenges collectively limit the efficiency and scalability of multimodal LLM training, resulting in MFU as low as ~20% in production-level training (§7.1).
To this end, we present MMScale, an efficient and adaptive framework to reform the training of multimodal LLMs. MMScale achieves state-of-the-art MFU which is close to unimodal (i.e., text-only) LLM training and effectively scales to large clusters with thousands of GPUs. The core of MMScale is adaptive resource allocation and data-aware reordering. These two techniques leverage the distinctive system characteristics of multimodal LLM training to eliminate model heterogeneity and data heterogeneity respectively.
For model heterogeneity, we meticulously analyze the pipeline bubbles stemming from model heterogeneity and identify their root causes (§2.2). Based on this analysis, we propose disaggregated training, which tailors the resource allocation and parallelism strategies across the modality encoder, LLM backbone, and modality generator. We also develop a novel and lightweight algorithm to navigate this complicated design space to choose the optimal resource and parallelism configurations. This innovative approach minimizes the pipeline bubbles and achieves optimal training efficiency.
For data heterogeneity, we categorize the training stragglers into inter-microbatch and intra-microbatch stragglers (§2.3) and introduce tailored algorithms that strategically reorder training data while maintaining synchronous training semantics. The inter-microbatch reordering algorithm reorders the data samples to evenly distribute the load across different data parallelism groups. The intra-microbatch reordering algorithm reorders the microbatches tailored for (interleaved) 1F1B pipeline scheme to minimize the pipeline bubbles. This two-level approach effectively mitigates training stragglers caused by data heterogeneity.
In addition, we customize system optimizations for multimodal LLM training to offload certain operations from the critical path (i.e., GPU computation). We implement an asynchronous data loader to pre-process the training data with negligible runtime overhead. We implement an in-house collective communication library, StepCCL, to hide the communication overhead within the computation.
In summary, we make the following contributions.
We present MMScale, an efficient and adaptive framework to reform the training of multimodal LLMs by harnessing the multimodal heterogeneity. It delivers state-of-the-art MFU on large-scale clusters with thousands of GPUs.
We identify and discuss the primary challenges associated with multimodal LLM training, which are summarized as model heterogeneity and data heterogeneity.
We implement disaggregated training across the three modules in multimodal LLMs. We propose adaptive resource allocation and data-aware reordering to effectively address model and data heterogeneity, respectively.
We implement MMScale and conduct experiments on our production cluster with thousands of GPUs. The experimental results show that MMScale achieves 54.7% MFU when training a 72B multimodal LLM on 1172 GPUs and outperforms Megatron-LM by up to 2.2 on throughput.
2 Motivation
2.1 Multimodal Large Language Model Training
Large language model. Large language models (LLMs) [18, 19, 20] have revolutionized natural language processing (NLP) by achieving state-of-the-art performance on a wide range of tasks, such as text generation, translation, and summarization. Many organizations have raced to develop their own LLMs, such as OpenAI’s GPT-4 [21], Google’s Gemini [11], and Meta’s Llama [22]. The core architecture of LLMs consists of a stack of homogeneous transformer layers [18] that use self-attention mechanisms to capture contextual information in text. LLMs are pre-trained with unsupervised learning on large-scale text corpora and then fine-tuned on task-specific text datasets. The text data is typically tokenized into fixed-length sequences, which are then fed into the model to learn the underlying patterns. The training process involves weeks or even months of computation on dedicated AI clusters with thousands of GPUs. According to one technical report [14], training 175B GPT-3 requires 4.6 million dollars. Optimizing the training process is essential to reduce the stupendous cost and accelerate the model deployment.
Multimodal LLM. The unimodal LLMs are limited to processing text data, which restricts their applicability to multimodal tasks (e.g., image understanding and generation). As a result, multimodal LLMs have emerged to address this limitation by integrating multiple modalities (e.g., images, audios, and videos) into the advanced LLMs [15], which support multimodal inputs and outputs during LLM generation. For example, GPT-4o [10] garners widespread attention by facilitating more natural interactions with humans through both visual and auditory modalities. Moreover, the predominantly text-based data in human societies is finite [23]. Harnessing multimodal data is inevitable to continually expand and enhance LLM capabilities.
Figure 1 illustrates the model architecture of a multimodal LLM, which consists of three modules: a modality encoder, an LLM backbone, and a modality generator [24, 15, 16]. The modality encoder transforms input data from different modalities (e.g., ViT [25] for images and Beats [26] for audios) into an intermediate representation (i.e., an embedding tensor), which is then projected into a unified embedding space across modalities with input projection layers (e.g., MLP and cross-attention). The LLM backbone, typically a transformer model (e.g., GPT [19, 20] and Llama [22]), processes the multimodal embeddings to discern the intricate data patterns and inter-modal relationships. The output data of the LLM backbone is subsequently refined by output projection layers, which tailor the information for each modality. Finally, the modality generator (e.g., Diffusion [27] for images and AudioLDM [28] for audios) transforms the LLM-processed information back into the respective modal outputs to generate coherent and contextually relevant results.
Multimodal LLM training necessitates training all three modules simultaneously. Additionally, during different training phases, specific modules are frozen to stabilize training loss and enhance model effectiveness [15]. Regarding the training data, the input sequence comprises text, image, and audio tokens. These data from different modalities are tokenized into subsequences which are then interleaved to form fixed-length training sequences [12]. In cases where the input sequence falls short, it is padded with zero tokens for batching. Distinct modality sub-sequences are demarcated by special tokens and processed through modality-specific encoders.
Training framework. To train the multimodal large language model with large-scale clusters, the de facto solution is to leverage Megatron-LM, a highly efficient and robust training framework for large-scale transformer models. Megatron-LM employs a unified parallelism strategy for the entire model. It combines tensor parallelism (TP) and pipeline parallelism (PP) to distribute the model parameters across multiple GPUs. TP divides the model parameter within each layer, while PP partitions parameters between layers. It also leverages data parallelism (DP) to distribute the training data. For multimodal LLMs, Megatron-LM is easily extended to integrate additional multimodal modules. Figure 2 illustrates the training topology of multimodal LLMs with Megatron-LM. Specifically, Megatron-LM treats the multimodal modules as additional layers within the LLM and incorporates additional PP stages to accommodate the modality encoder and generator. The same TP strategy used in the LLM backbone is applied to these two multimodal modules. If the modality encoder and generator are not large enough, they are replicated across the GPUs in the TP group to maximize resource utilization. As for DP, Megatron-LM applies the same DP strategy to the multimodal modules as the LLM backbone. The projector is co-located with the modality encoder and generator and is replicated across the GPUs in the TP group.
However, this training framework introduces significant computation imbalance stemming from model heterogeneity and data heterogeneity due to its rigid resource allocation method (i.e., the multimodal modules share the same DP and TP strategies with LLM backbone). It only achieves ~20% MFU (§7.1) which is significantly lower than the MFU of ~50% observed in training unimodal (text-only) LLMs [29].
2.2 Model Heterogeneity
(a) Distribution of text subsequence size. (b) Distribution of image subsequence size. (c) Distribution of image subsequence count.
The first challenge is the computation imbalance arising from model heterogeneity. Each module in multimodal LLMs bears different computational demands due to varying operators and inputs. For instance, ViT, as modality encoder, is constructed with narrow transformer layers (i.e., small hidden size), whereas LLM backbone is built on broader transformer layers (i.e., large hidden size). Meanwhile, Diffusion, as modality generator, utilizes a combination of convolution and attention layers (i.e., U-Net). This architectural diversity results in distinct computation time for each module. Figure 3 shows varying forward time under different input configurations with Megatron-LM. We demonstrate one PP stage of LLM backbone with PP size of 10 and TP size of 8. The first configuration parameter is the number of images in the 8K input sequence, and the second is the image resolution. The time differs markedly across different configurations.
The computational imbalance between modules leads to two types of pipeline bubbles in pipeline parallelism. The first arises in the modality encoder and generator stages, as shown in Figure 4(a), resulting from their inadequate utilization of assigned GPU resources. The second type emerges in the stages of the LLM backbone, as shown in Figure 4(b). This is because the intensive computational demands of the encoder and generator extend their stage durations. Due to the pipeline dependency, the LLM stages are forced to wait for the multimodal stage to complete, thereby creating pipeline bubbles. The latter problem is particularly pronounced during large-scale multimodal LLM training, where the bulk of GPU resources are allocated to the LLM backbone. These pipeline bubbles, stemming from model heterogeneity, substantially diminish the MFU during the training.
2.3 Data Heterogeneity
The second challenge is the computational imbalance stemming from data heterogeneity. Each input sequence (i.e., training sample) for multimodal LLM training consists of interleaved modality subsequences that exhibit highly skewed distributions. Focusing on images and texts, we perform data characterization on the LAION-400M dataset [30], an open-source collection of images paired with text captions. Each image (i.e., one image subsequence) is segmented into 1616 patches, and each patch is tokenized into one image token. The texts are tokenized through Llama tokenizer. The image tokens are interleaved with text tokens to create an 8K-token input sequence for training. As shown in Figure 5(a) and Figure 5(b), the sizes of text and image subsequences display distinctly skewed distributions. We further analyze the count of modality subsequence per training sample using image as an example. The count of image subsequences per training sample, shown in Figure 5(c), also demonstrates a skewed distribution. Different sample size (i.e., modality tokens per sample) leads to varying computation time in the modality encoder and generator stages.
Such data heterogeneity results in both intra-microbatch and inter-microbatch stragglers within the pipeline parallelism (PP) stages of the modality encoder and generator. These stragglers exacerbate the computational imbalances and further reduce the GPU utilization. It is noted that all microbatches within LLM backbone have the same computation time since the sequence length is fixed. We do not consider data heterogeneity between global batches, as each global batch contains numerous randomly shuffled training samples (e.g., thousands with a large DP size), which effectively smooths out the data heterogeneity.
Intra-microbatch straggler. Intra-microbatch straggler occurs when particularly large training samples decelerate training, as DP groups handle variably-sized training samples. Illustrated in Figure 6, the first DP group () processes two large training samples within two microbatches. In contrast, the second DP group () processes two smaller samples in the same microbatches, completing them more swiftly. Consequently, lags behind and becomes the straggler, which delays the overall training process.
Inter-microbatch straggler. Inter-microbatch straggler emerges from pipeline imbalances between microbatches. As depicted in Figure 7, the first pipeline stage is the modality encoder followed by one LLM backbone stage. Figure 7(a) illustrates the pipeline without data heterogeneity, where the modality encoder processes each microbatch with consistent time. In contrast, Figure 7(b) depicts the pipeline with data heterogeneity, where the forward time of the modality encoder varies markedly across microbatches. The straggler (i.e., the microbatch ) significantly delays the training process of the subsequent PP stages, leading to a large pipeline bubble.
3 MMScale Overview
We present MMScale, an efficient and adaptive framework to reform the training of multimodal LLMs by harnessing the multimodal heterogeneity. MMScale exploits the characteristics of parallelism training of multimodal LLMs to achieve high efficiency and scalability. MMScale eliminates the model heterogeneity by adaptive resource allocation (§4) and harnesses the data heterogeneity by data-aware reordering (§5). In addition, MMScale adopts some system optimizations tailored for multimodal LLM training (§6). Here we provide a brief overview of MMScale as Figure 8 shows.
MMScale orchestrator. Before training, MMScale employs a training orchestrator to determine the resource allocation and parallelism strategy for each module in multimodal LLMs. This scheduler first gathers the model architecture and training configuration (e.g., global batch size) from the user and randomly samples a subset of training data to analyze the data distribution. Utilizing this information, it runs a series of benchmarking training trials and constructs a performance profiler with linear interpolation to estimate each module’s computation and communication time. Based on the profiling results, the training orchestrator decides the optimal resource allocation and parallelism strategy with adaptive resource allocation for one specific training task, as detailed in §4.
MMScale initializer. Then, MMScale initializes the parallelism units for the modality encoder, LLM backbone, and modality generator, respectively. MMScale allocates different numbers of GPUs to each unit. Each unit then establishes the communication group across different nodes and GPUs. The unit loads the model checkpoint from distributed file system and shards the model parameters and optimization states. Finally, MMScale conducts several communication trials to warm up the system and test connectivity.
MMScale runtime. After initialization, MMScale begins the main training process. In each training iteration, it first retrieves a global batch of training samples from the distributed file system for pre-processing. Then, MMScale performs data-aware reordering to reorder the training samples within one global batch without breaking the synchronous training semantics [31]. This module effectively eliminates both inter-microbatch and intra-microbatch data heterogeneity, as detailed in §5. The data then undergoes sequentially through the modality encoder, LLM backbone, and modality generator in the training pipeline. Finally, it synchronizes the gradients and model parameters through the all-gather operation, employing the ZERO-1 optimization [32] and mixed precision training [33]. Additionally, MMScale adopts a dedicated process to periodically and asynchronously save the model checkpoint to the distributed file system for fault tolerance.
4 Addressing Model Heterogeneity
Before diving into the details, we first introduce the disaggregation of the training across the three modules (i.e., modality encoder, LLM backbone, and modality generator) in multimodal LLMs. This disaggregation facilitates adaptive GPU resource allocation among the different modules based on the training workload. Subsequently, we formulate the problem of adaptive resource allocation to minimize the training time per iteration. Finally, we present the detailed algorithm to optimally address the problem caused by model heterogeneity.
4.1 Disaggregated Training
Disaggregating the training among the three modules is the foundation of adaptive resource allocation in MMScale. Figure 9 demonstrates the training topology of disaggregated training in MMScale. Different from the rigid resource allocation in Megatron-LM (i.e., Figure 2), MMScale is able to adaptively adjust the resource allocation. For instance, MMScale allocates 4 GPUs (DP=2 and TP=2) to the modality encoder, 12 GPUs (DP=3 and TP=4) to the LLM backbone per PP stage, and 4 GPUs (DP=1 and TP=4) to the modality generator. Additionally, the projector layers are colocated with either the modality encoder or generator, with their number of replicas adapting as needed. We implement the disaggregated training through a dedicated module, i.e., parallelism unit.
Parallelism unit. At training initialization, we need to establish the communication group according to the resource allocation and parallelism strategy. MMScale introduces a module, parallelism unit, composed of one or more PP stages. Each unit can adopt its own DP and TP strategies and form a specific communication group. Inter-unit connections are facilitated by a communication broker, which bridges PP communication across parallelism units. Users are only required to specify the DP and TP configurations for each parallelism unit, and MMScale automatically sets up the communication group and communication broker. MMScale treats the modality encoder, LLM backbone, and modality generator as three individual parallelism units. The detailed implementation of parallelism unit is discussed in §6.
4.2 Problem Formulation
With disaggregated training, we are able to adaptively allocate GPU resources across the three modules. The problem now lies in determining the optimal resource allocation and parallelism strategy to minimize the training time per iteration. Exhaust search is infeasible due to the large search space of the parallelism strategy and cluster size. One strawman solution is to allocate the GPU resources proportional to the model flops of each module. However, this method falls short as it overlooks complex patterns in parallelism training. For instance, different DP and TP sizes in modality encoder could lead to different computation times while the model flops are the same. Before diving into the adaptive resource allocation algorithm, we first formulate the optimization problem.
LLM backbone. We begin by formulating the LLM backbone, focusing on the forward pass as the backward pass time mirrors this. In LLM training, microbatch size is set to one to prevent GPU memory overflow. Assume the global batch size for one iteration as and the TP size of the LLM backbone as . Let the PP and DP size of LLM backbone be and . The number of GPUs allocated to LLM backbone is . Let the forward time (including communication time) of the entire LLM be , where represents forward time function. Therefore, the forward time of one PP stage for one microbatch is . Besides, the number of microbatches per iteration is .
Modality encoder and generator. In MMScale, the modality encoder is regarded as a parallelism unit with PP size . Let the TP size be and the DP size be . The number of GPUs allocated to modality encoder is . The microbatch size is which is determined by the LLM backbone. Let the forward time (including communication time) of the entire modality encoder be . The forward time of one PP stage for one microbatch in the modality encoder is . Similarly, the forward time of one PP stage in the modality generator is , where is the number of GPUs allocated to modality encoder.
Objective function. Based on the preceding analysis, we next define the objective function for the optimization problem, i.e., the training time of one iteration. As shown in Figure 10, we demonstrate the pipeline of forward pass in multimodal LLM training. The LLM backbone comprises two PP stages, whereas the modality encoder and generator each consist of one PP stage. The training process is categorized into two phases: warm-up phase and steady phase. The warm-up phase spans from the initiation to the completion of the first microbatch to populate the pipeline. This phase’s duration is calculated as , which is formulated as follows.
The steady phase’s duration is dominated by the maximal computation time among PP stages, which is calculated as , where is the number of microbatches per iteration. It is formulated as:
Therefore, the objective function is to minimize . For the backward pass, the objective function remains analogous to that of the forward pass. Adjustments are made by changing , and from forward time functions to the sum functions of forward and backward time. This formulation holds for GPipe and 1F1B. We will retrofit the formulation to adapt to VPP later. As for the communication time, it consists of the time of P2P communication in PP and collective communication in DP. It is modeled as the communication volume divided by the communication bandwidth. If the sizes of TP and DP are fixed, the communication time remains constant.
Constraints. Besides the objective function, we must consider constraints to ensure training feasibility. The first constraint is the resource constraint. The number of GPUs allocated to each module should be where is the total number of GPUs in the cluster. The second constraint involves memory. We consider LLM backbone. Memory allocation involves four parts: model parameters, gradients, optimizer states, and activation states. The memory of model parameters and gradients on one GPU is calculated as: , where denotes total memory for the LLM parameters and gradients. The memory for optimizer states on one GPU (with ZeRO-1 optimization) is: , where denotes the total memory for the optimizer states. ZeRO-1 partitions the optimizer states across DP groups. The peak memory for activation states on one GPU is: , with representing the memory needed for one microbatch of activation states across the entire LLM. In 1F1B, the first PP stage requires storage for microbatches of activation states. We eschew using GPipe in MMScale since GPipe consumes more memory. The memory constraint ensures the sum of the four memory parts on one GPU doesn’t exceed GPU capacity. As for modality encoder and generator, the formulation is similar.
4.3 Adaptive Resource Allocation
The optimization problem is non-convex, with , and the DP, TP sizes as positive variables. is predefined by the user. Solving this with an exhaust search algorithm to explore all possible variable values is impractical due to the extensive search space, particularly in large clusters. Designing an efficient algorithm that quickly identifies the optimal resource allocation and parallelism strategy is a significant challenge.
Convex optimization on resource allocation. Our key insight is to decompose the non-convex optimization problem into a series of simplified convex problems with variables (i.e., the resource allocation). We confine the TP size to on an NVIDIA GPU node with 8 GPUs and adjust the DP size as a factor of to balance the computation across DP groups. The PP size of the LLM backbone is calculated as . The set of possible parallelism strategies is a manageable and finite set, i.e., the Cartesian product of the TP and DP size. This allows us to enumerate all feasible TP and DP sizes and transform the original optimization problem into a set of simplified problems. In the simplified problem, the objective function is maximal and additional of the functions: , and , where are positive. Therefore, the objective function is convex. Similarly, the constraint functions are also convex. As a result, the simplified optimization problem is convex and can be efficiently solved to optimality by existing solvers [34, 35]. The algorithm, detailed in Algorithm 1, efficiently finds the optimal resource allocation and parallelism strategy to minimize iteration time.
Virtual pipeline parallelism (i.e., interleaved 1F1B) reduces the warm-up time by dividing model into finer-grained virtual PP (VPP) stages. Each PP stage contains VPP-size VPP stages. In the warm-up phase, each PP stage launches the computation of one VPP stage, and the warm-up time is divided by VPP-size. To align our formulation with VPP, we proportionally reduce the warm-up time based on VPP-size.
5 Addressing Data Heterogeneity
The data heterogeneity causes intra-microbatch and inter-microbatch stragglers. Addressing both straggler simultaneously renders the problem intractable due to its high complexity (e.g., the order of the training samples results in factorial-level complexity) and the necessity of runtime execution. To tackle this challenge, we decompose the overarching problem into two tractable subproblems. Initially, we present intra-microbatch reordering to eliminate stragglers across DP groups. We then introduce inter-microbatch reordering to minimize pipeline bubbles caused by the straggler microbatches. The combination of the two reordering, i.e., data-aware reordering, effectively addresses the data heterogeneity.
5.1 Intra-microbatch Reordering
Insight. To address the first subproblem of intra-microbatch stragglers, we identify the straggler by pinpointing the DP group with the largest training samples. As illustrated in Figure 6, the first DP group becomes a straggler as it contains the two largest training samples. To neutralize this imbalance, we propose reordering the training samples within the global batch by size. Specifically, as depicted in Figure 11, we reorder the training samples into the sequence . This strategy effectively distributes the computational load more evenly and improves the scalability. It eliminates the straggler without breaking the synchronous training semantics [31], as the reordering occurs within an individual global batch.
Intra-microbatch Reordering. Leveraging this insight, we propose intra-microbatch reordering to balance computational load and improve the overall scalability. Formally, the challenge involves minimizing the maximum computation time among DP groups. This problem corresponds to the NP-hard problem, multiway number partitioning [36], which aims to minimize the largest sum. There is no known polynomial-time algorithm capable of determining the optimal solution. Given the variability of runtime training samples, any algorithm employed must run before each iteration with negligible runtime overhead. Consequently, we adopt the greedy number partitioning algorithm, which guarantees an approximation ratio of [37]. The detailed algorithm is summarized in Algorithm 2. The function receives the original training samples and DP size . This algorithm first sorts the training samples in ascending order by the sample size (line 3). Then, it loops over the training samples and assigns the sample to the DP group with the current lowest computational load (line 6-8). It then returns the reordered training samples (line 9-11). The algorithm is lightweight and maintains a time complexity of .
5.2 Inter-microbatch Reordering
The second subproblem is the inter-microbatch straggler. As we discussed in §2.3, data heterogeneity leads to varied computation times across microbatches within the modality encoder and generator. The straggler microbatch prolongs the training by creating large pipeline bubbles. In the context of 1F1B pipeline scheme, the overall iteration time is primarily governed by pipeline bubbles and the computation time at the first PP stage of the modality encoder, as illustrated in Figure 12. Let the PP size be and the number of microbatches be ( and in Figure 12). Typically, is larger than to reduce the proportion of time spent in the warm-up and cool-down phases. We abstract an innovative concept, pipeline intervals, at the first PP stage. As shown in Figure 12, these intervals are typically filled with forward pass, except for the last intervals (i.e., , , and ). Straggler microbatches in either the encoder or generator prolong these intervals or increase the unfilled area (i.e., bubble). Designing an efficient runtime algorithm to minimize pipeline bubbles presents a significant challenge.
Insights. We leverage two insights to solve this problem. The first insight involves minimizing the volume of intervals that cannot be filled. As shown in Figure 12, the last intervals (i.e., to ) remain unfilled. These intervals become the pipeline bubbles and prolong the training iteration. We observe a positive correlation between the volume of and the size of the microbatch. For instance, is significantly larger than and . By strategically reordering the training samples to position the smallest microbatches at the end, we are able to reduce these unfilled intervals (i.e., pipeline bubbles).
The second insight involves minimizing the unfilled area of front intervals. As shown in Figure 12, the first interval is filled with by the to microbatches’ forward passes. The volume of the first interval (i.e., ) is determined by the aggregate forward and backward times of the first microbatch in LLM backbone and modality generator. Such characteristic allows the accommodation of large microbatches early in the sequence to fill the first interval. For subsequent intervals, is filled by the microbatch’s forward pass. By evaluating the volume of , we place the microbatch, whose forward time most closely matches this volume, at the position, thereby reducing the unfilled area (i.e., pipeline bubbles).
Inter-microbatch Reordering. Based on the two insights, we design runtime inter-microbatch reordering to minimize the pipeline bubbles. This algorithm is designed for 1F1B pipeline scheme. We will retrofit the algorithm to VPP (i.e., interleaved 1F1B) later. Algorithm 3 summarizes the pseudo code. The function receives the original order of microbatches and the PP size . It stores reordered microbatches in and pending microbatches in (line 2). Early in the process, the smallest microbatch is placed at first to activate all pipeline stages promptly (line 3). Subsequently, it selects the smallest microbatches and places them at the end to minimize unfilled intervals (line 4 and line 12). It then loops over the remaining microbatches (line 5-11). In each loop iteration, it calculates the volume of the interval through the function . For the first interval, it selects microbatches that closely match the interval volume in sum forward time; for others, it selects a single microbatch whose forward time aligns closely with the interval volume. This loop ensures maximal filling of the remaining intervals, which minimizes pipeline bubbles.
The functions and operate with a time complexity of . The function calculates interval volumes using the current order . This calculation is facilitated by a dynamic programming algorithm that utilizes a recursive formula derived from pipeline dependencies. Specifically, each microbatch’s start time depends on two factors: the completion of the preceding microbatch on the same device and the availability of input data from the upstream microbatch. Consequently, the end time of each microbatch is determined by the maximum of these two dependencies plus its own computation time. This dynamic programming algorithm exhibits a complexity of per function invocation. The algorithm is lightweight and efficient with a time complexity of .
Virtual pipeline parallelism (i.e., interleaved 1F1B) also follows the one forward and one backward pipeline scheme to reduce the memory footprint. The fundamental insights of our algorithm apply universally to any 1F1B-based pipeline, including VPP. We adapt the algorithm by computing multiple (i.e., VPP size) intervals and filling them with the corresponding number of forward passes from a single microbatch.
6 System Implementation and Optimization
We implement MMScale with 6.3K lines of code in Python and C++, and integrate it with Megatron-LM [17], a state-of-the-art training framework for large language models. MMScale is able to support multimodal LLM training with a variety of modalities on a large-scale GPU cluster (e.g., thousands of GPUs used in §7). MMScale leverages a distributed file system to store the training data and model checkpoints. It handles failures by automatically recovering the training process from the latest checkpoint.
MMScale orchestrator. MMScale’s training orchestrator, implemented as a Python script on a dedicated CPU node, formulates the adaptive resource allocation problem using Disciplined Convex Programming [38]. It employs the CVX solver [39] to efficiently solve this problem within milliseconds. The orchestrator then records the optimal resource allocation and parallelism strategy to a configuration file, which the Kubernetes controller uses to launch the training task.
MMScale initializer. MMScale incorporates parallelism unit that manages specific parallelism strategies for each module in multimodal LLMs. Parallelism unit is initialized using PyTorch Distributed [40] library with NCCL as the default communication backend, except for the TP communication where we use StepCCL instead. MMScale shards the model parameters in accordance with the established parallelism strategy.
Parallelism unit. As we discussed in §4.1, the disaggregated training is implemented through a specific module: parallelism unit. When initializing the distributed training, MMScale first establishes the communication groups within one parallelism unit. Each GPU process has a global rank and a local rank within the unit, which facilitates the distributed initialization. Then, it initializes the communication broker to establish the PP communication between adjacent parallelism units. All communication traffic between parallelism units is routed via the communication broker.
We implement the communication broker by modifying the batched send and receive operations in Megatron-LM to separate operations. This allows flexible communication between multiple upstream and downstream GPU processes. The communication broker adjusts the communication data by concentrating and scattering the data as needed while maintaining the data order. Strategically located on the GPU of the last PP stage in the upstream unit or the first PP stage in the downstream unit, the communication broker avoids additional communication overhead. Besides, the number of communication brokers between two units is determined by the greatest common divisor of the two parallelism units’ DP sizes to maximize the communication bandwidth. Moreover, in the Megatron-LM, the reliance on synchronous communication compels upstream stages to pause until downstream stages fully receive the data. This introduces unnecessary dependencies in the pipeline. To alleviate this, we implement asynchronous send operations that eliminate these superfluous dependencies and redesign the communication topology to prevent potential deadlocks.
Asynchronous data loader. When training multimodal LLMs, training samples often combine lightweight text with heavyweight multimodal data. The latter significantly increases data loading time. For example, a typical training sample could include a 256-word text sequence and ten 10241024 RGB images. The text is just kilobytes, whereas the images are total of 120 megabytes. Loading multiple such samples can take several seconds. MMScale’s data loader employs a producer-consumer model to enhance efficiency. The producer, operating on dedicated CPU nodes, fetches data from the distributed file system and pre-processes training data (e.g., image cropping) asynchronously with the GPU training process. The consumer, integrated within the main training process, receives this pre-processed data and reorders it to eliminate the data heterogeneity. The producer and consumer communicate through RPC calls, and use RDMA network for lower latency if available.
Mitigating TP overhead with StepCCL. Tensor Parallelism (TP) is commonly adopted to facilitate training large Transformer-based models with multiple GPUs connected with high bandwidth (e.g., with NVLinks). Specifically, TP divides the linear layers, into smaller sub-modules, which are then distributed across the GPUs. After parallel computation, all GPUs perform collective communication to aggregate data and produce identical results. The TP communication overhead severely degrades overall performance, especially on A800 and H800 GPUs with restricted NVLink bandwidth.
We implement the communication overlap with an in-house collective communication library called StepCCL to reduce the TP overhead. StepCCL is a PyTorch custom plugin that performs cross-GPU collective communication, including allgather, reduce-scatter, and allreduce, which is similar to NCCL. However, NCCL occupies several CUDA Streaming Multiprocessors (SMs) for executing its communication kernel and is known to harm the performance of its concurrent GEMM (matrix multiplication) [41]. To solve this, StepCCL leverages the DMA engine directly to perform data transmission without using any SM at all. This enables StepCCL and GEMM to run simultaneously on a GPU without slowing down each other. This cornerstone facilitates our subsequent development of communication overlap.
Figure 13 shows an example of how StepCCL works in overlapping the allgather (AG) operation with GEMM. We start by decomposing the single GEMM and the corresponding communication into several smaller pairs. Each small communication operation starts sequentially on a communication stream, with its paired GEMM executed on the default computation stream. The communication overhead is fully hidden except for the first allgather.111If the number of allgather/GEMM is large enough, the only allgather in the critical path should have negligible overhead. But dividing a large GEMM into finer granularity sometimes could lead to overall slowdown. In practice, the number is actually configurable. After all GEMMs finish, we perform an extra layout remapping operation (usually with negligible overhead) to ensure identical results with the baseline. Figure 14 describes the details of the layout remap process. In some rare cases during backward propagation, we find the remap overhead is high due to certain model dimension. To mitigate this, we further overlap the remap with the computation of the weight gradients, so eventually we nearly get the full performance gain of the communication overlap.
Finally, although the overlap idea is also studied in many related works [42, 43, 44], we highlight the key differences of ours. Unlike prior work that fuses GEMM with TP communication into one CUDA kernel [42, 43], we choose a modular design and do not use fusion for more flexibility. For example, when TP communication is longer than GEMM, fusing them cannot fully hide the communication overhead. However, with the modular design, we are able to hide the communication with other modules without dependency (e.g., in cross-attention), which is not possible with the fused implementation. This enables broader adoption of StepCCL in many other scenarios.
7 Evaluation
In this section, we first use large-scale experiments to demonstrate the overall performance improvements of MMScale over existing multimodal LLM training solution. Next, we use microbenchmarks to deep dive into MMScale and show the effectiveness of each component of MMScale. Finally, we conduct a case study to evaluate MMScale under different frozen settings and evaluate the system optimizations.
| Models | # of | Hidden | FFN Hidden | # of | # of |
| Layers | Size | Size | Heads | Groups | |
| Llama3-7B | 32 | 4096 | 11008 | 32 | 32 |
| Llama3-13B | 40 | 5120 | 13824 | 40 | 40 |
| Llama3-70B | 80 | 8192 | 28672 | 64 | 8 |
Setup. Our experiments are conducted on a production GPU cluster for multimodal LLM training, with each node equipped with eight NVIDIA A800 GPUs, 1TB of memory, and 96 vCPUs. GPUs within one node are interconnected by 300GB/s (bidirectional) NVLink, while nodes are connected by 4*200 Gbps RDMA network based on RoCEv2 with rail-optimized topology. The overall experiments use up to 1296 GPUs, and the microbenchmark utilizes up to 98 GPUs. We use PyTorch 2.1.2 and NVIDIA CUDA 12.2 for our evaluation.
Models. For LLM backbone, we choose the representative LLM architecture, Llama3 [22], which is widely used in both academia and industry. Table 1 lists the detailed model configurations. As for the modality, we focus on images and texts. MMScale is also compatible with other modalities. For modality encoder and modality generator, we use ViT-Huge [45] (0.63B) and Stable-Diffusion 2.1 [46] (1B) respectively. These two models are widely used for image understanding and generation. The three LLM backbones (i.e., Llama3-7B, Llama3-13B, and Llama3-70B) are paired with ViT-Huge and Stable-Diffusion to form multimodal LLMs designated as MLLM-9B, MLLM-15B, and MLLM-72B. For large multimodal LLM (i.e., MLLM-72B), we use high image resolution (i.e., 10241024) for generation since the large LLM is able to process more context information. For small models, we use low image resolution (i.e., 512512).
Datasets. For our experiments, we use the representative open-source dataset, LAION-400M. We generate training data by interleaving the image and text subsequences, forming input sequences up to 8192 tokens long. This dataset is also employed in our production multimodal LLM training. As detailed in §2.3, each training sample includes a varying number of image tokens and text tokens, which introduces data heterogeneity in multimodal LLM training.
Metrics. We use the Model FLOPs Utilization (MFU) as the primary metric to evaluate MMScale. MFU measures the percentage of GPU FLOPs that are effectively utilized during model training. We also use the training throughput (TPT.) to evaluate the training speed of MMScale. Since MMScale and Megatron-LM may utilize different numbers of GPUs due to varying resource allocation strategies, we also indicate the number of GPUs used in the throughput charts.
7.1 Overall Performance
Setup. We first compare the overall performance of MMScale against Megatron-LM on a large-scale GPU cluster (up to 1296 GPUs). We retrofit Megatron-LM to support multimodal LLM training by integrating modality encoder and generator into the training pipeline. Megatron-LM employs rigid resource allocation as described in §2.1. In Megatron-LM, we set the PP size of the LLM backbone to 1, 2, and 10 for Llama3-7B, Llama3-13B, and Llama3-70B, respectively. TP size is set to 8. As for MMScale, the parallelism strategy is determined by adaptive resource allocation. In our experiments, one GPU is able to facilitate training ViT and Stable-Diffusion. For both Megatron-LM and MMScale, we replicate the encoder and generator across the GPUs within the TP group to process different image subsequences, whereas TP itself is not used. We set global batch size to 1920.
The experimental results are shown in Figure 15 and Figure 16. Figure 15 shows the MFU of MMScale and Megatron-LM under different multimodal LLMs. Figure 16 shows the training throughput and marks the number of GPUs used in each experiment. We summarize the experiment as follows.
As shown in Figure 15, MMScale achieves 51.8%-54.7% MFU under large-scale multimodal LLM training scenario. This performance closely approximates that of state-of-the-art unimodal (i.e., text) LLM training [29], which demonstrates the effectiveness of MMScale in eliminating the model and data heterogeneity in multimodal LLM training.
MMScale significantly outperforms Megatron-LM, delivering 1.7-2.8 the MFU and 1.7-2.2 the training throughput when training MLLM-9B and MLLM-15B with a similar number of GPUs. These performance gains largely stem from MMScale’s adaptive resource allocation strategy. Megatron-LM’s rigid strategy often leads to GPU underutilization in multimodal modules, since it assigns too many GPUs to the modality encoder and generator. In contrast, MMScale adaptively adjusts GPU allocations based on specific model and data demands. Additionally, MMScale’s data-aware reordering technique further boosts efficiency.
In MLLM-72B training scenario, MMScale also outperforms Megatron-LM by 1.2 on MFU and 1.3 on training throughput with a similar number of GPUs. The high image resolution prolongs the execution time of the multimodal module, which introduces pipeline bubbles in LLM backbone. MMScale addresses this by allocating additional GPUs to these modules to balance the pipeline. MMScale’s data-aware reordering strategy continues to diminish data heterogeneity, thereby increasing overall training efficiency.
7.2 Deep Dive into MMScale
In this subsection, we perform microbenchmarks to evaluate the effectiveness of each MMScale’s component. Our experiments utilize up to 98 NVIDIA A800 GPUs. We set the global batch size to 128, 64, and 40 for MLLM-9B, MLLM-15B, and MLLM-72B, respectively.
7.2.1 Adaptive Resource Allocation
We measure the MFU and training throughput of MMScale and other resource allocation strategies, including a comparison with Megatron-LM and a ratio-based approach. The ratio-based strategy allocates GPUs according to the computational demands (flops) of each module. We also annotate the total number of GPUs utilized in the throughput chart. The experimental results are shown in Figure 17. MMScale consistently outperforms the baseline strategies, achieving 1.3%-2.7% higher MFU and 1.4%-2.7% higher training throughput. Although the ratio-based strategy outperforms Megatron-LM’s rigid allocation, it still lags behind MMScale since it neglects the intricate performance model (§4.2) of multimodal LLM training. MMScale employs an adaptive resource allocation strategy that optimally balances computational loads across the three modules and achieves high resource utilization.
We also evaluate the running time of MMScale’s resource allocation algorithm under different training settings, as detailed in Table 2. The algorithm completes the allocation process in under one second. The overhead is negligible compared to the days or even weeks required for overall training.
7.2.2 Data-aware Reordering
We evaluate the effectiveness of MMScale’s data-aware reordering by comparing it against the random order set by PyTorch’s DataLoader, while keeping other components the same. The effectiveness is gauged through metrics such as MFU and training throughput (TPT.). We use the optimal resource allocation and parallelism strategy decided by MMScale’s adaptive resource allocation algorithm. Given that the GPU allocation policy remains unchanged, the number of GPUs is not specified in this experiment. The experimental settings are the same as those in §7.2.1. The results are shown in Figure 18. MMScale consistently outperforms the baseline, achieving 1.03%-1.11% higher MFU and training throughput. The performance gap becomes more pronounced as the model size decreases. This is because the smaller model size leads to a higher data parallelism (DP) size, which causes more inter-microbatch heterogeneity. In essence, MMScale’s data-aware reordering effectively mitigates data heterogeneity and enhances the training efficiency.
(a) MFU. (b) Throughput.
| Model | # of | Global | Algorithm |
| GPUs | Batch Size | Overhead | |
| MLLM-72B | 1296 | 1920 | 922ms |
| MLLM-72B | 648 | 960 | 641ms |
| MLLM-72B | 324 | 480 | 441ms |
| MLLM-72B | 112 | 240 | 133ms |
MMScale’s data-aware reordering algorithm executes before each training iteration. Table 3 presents a detailed analysis of the runtime overhead of the reordering algorithm. The DP size is the number of data parallelism groups in modality encoder. It completes the reordering in just a few milliseconds, a duration that is negligible compared to the several seconds to tens of seconds required for one iteration.
7.3 Case Study
In this subsection, we first evaluate MMScale under different frozen training settings. We then evaluate the effectiveness of system optimizations tailored for multimodal LLM training.
7.3.1 Frozen Training
We conduct a frozen training experiment under four specific training settings: complete module freezing, exclusive encoder training, exclusive LLM training, and exclusive generator training. We keep training the projectors. In these scenarios, frozen modules neither compute weight gradients nor update weights. All other experimental setup aligns with those detailed in §7.2. We evaluate the MFU and training throughput (TPT.) of MMScale compared to Megatron-LM. The experimental results are presented in Figure 19 and Figure 20. MMScale consistently outperforms Megatron-LM across all frozen training configurations, achieving 1.4%-2.9% higher MFU and 1.2%-2.9% increased training throughput. This pronounced performance gap underscores the challenges posed by Megatron-LM’s rigid resource allocation in complex training environments. In contrast, MMScale adaptively adjusts resource allocation based on training settings and consistently achieves high resource utilization.
(a) MFU. (b) Throughput.
| Model | DP | Global | Intra-microbatch | Inter-microbatch |
| Size | Batch Size | Reordering | Reordering | |
| MLLM-72B | 30 | 1920 | 1.82ms | 5.87ms |
| MLLM-72B | 60 | 1920 | 1.87ms | 3.02ms |
| MLLM-72B | 120 | 1920 | 1.89ms | 1.71ms |
(a) All modules freezing. (b) Exclusive encoder training. (c) Exclusive LLM training. (d) Exclusive generator training.
(a) All modules freezing. (b) Exclusive encoder training. (c) Exclusive LLM training. (d) Exclusive generator training.
7.3.2 System Optimizations
Asynchronous Data Loader. We conduct an experiment to evaluate the effectiveness of asynchronous data loading optimization. Setting the DP size to one, we measure the average data loading time per iteration solely on the CPU. We then compare data loading times with and without this optimization and use varying numbers of images and different image resolutions for one training iteration. The results, depicted in Figure 21, indicate that asynchronous data loading significantly reduces loading time from seconds to milliseconds. The first parameter in the x-axis represents the number of images, while the second parameter denotes the image resolution. This reduction renders the loading time negligible relative to the total training time.
Mitigating TP overhead with StepCCL. To evaluate the effectiveness of StepCCL in mitigating the TP overhead, we conduct an experiment that measures the iteration time of the LLM backbone with training of one single PP stage (i.e., one minimal TP group) under various TP sizes. We compare the iteration time with and without StepCCL enabled. The results are shown in Figure 22. StepCCL significantly reduces the iteration time by overlapping the TP communication with computation. It outperforms the baseline by 1.1-1.17 when the TP size is 4 and 1.15-1.17 when the TP size is 8. The gains are more pronounced at large TP size, where communication overhead is more substantial. These findings confirm that StepCCL effectively mitigates TP overhead.
8 Discussion
Parallelism optimization. Many studies [47, 48, 49] refine parallelism strategies for deep learning models, but they fall short for multimodal LLMs due to their expansive search spaces and large cluster requirements. These methods separately optimize different strategies, resulting in inefficient parallelism Additionally, these methods generally assume data homogeneity, standardizing computations across all training samples. This assumption does not hold for multimodal LLMs due to the data heterogeneity in modality inputs of each sample. In contrast, MMScale leverages the specific training pattern of multimodal LLMs, creating a customized resource allocation problem that integrates tensor, pipeline, and data parallelism simultaneously. Besides, MMScale leverages data-aware reordering to harness the data heterogeneity.
Sequence and expert parallelism. Sequence parallelism (SP) [50] is designed to partition the training sequence into multiple subsequences for parallel training. It addresses the challenges of processing long sequences in LLMs. Expert parallelism (EP) [51], specifically devised for mixture-of-experts (MoE) LLMs [52], enables parallel training of multiple feedforward network (FFN) experts. These parallelism strategies are orthogonal to multimodal LLM training. In MMScale, both SP and EP are integrated into the training framework. MMScale treats SP and EP sizes as predefined parameters in the adaptive resource allocation optimization problem.
Disaggregating multimodal modules from LLM. In MMScale, we address model heterogeneity by treating multimodal modules and LLM as separate stages in pipeline parallelism (PP). Another approach might integrate some LLM layers with multimodal modules into the same PP stage, while distributing the remaining LLM layers to maintain balance. However, this often becomes suboptimal since it requires multimodal modules to adopt the same TP strategy as the LLM layers within the same PP stage. Such alignment degrades performance when the optimal TP strategies for multimodal modules differ from those of LLM layers. Therefore, we choose to disaggregate multimodal modules from LLM as separate PP stages to achieve better performance.
(a) TP=4. (b) TP=8.
9 Related Work
LLM training. Many efforts have been made to optimize the training of LLMs from system perspectives. For LLM pretrain, Megatron-LM [17] and DeepSpeed-Megatron [53] propose customized 3D-parallelism and are de facto standards for training large LLMs. DeepSpeed-ZeRO [32] and Pytorch-FSDP [54] reduce redundant memory consumption in data parallelism. HPN [55] proposes a new dual-ToR network architecture trailed for LLM training to reduce ECMP conflict. A set of works [56, 29, 42, 43, 44] overlap the communication and computation operators in LLM training. Fault tolerance through replication and checkpoint is advanced in large training clusters by studies [29, 57]. Efforts like [58, 59, 60] further optimize recovery process in cloud spot instance scenarios. These system optimizations of LLM training are orthogonal to MMScale. They overlook the model and data heterogeneity in multimodal LLMs. MMScale also integrates several of these optimizations in training the LLM backbone.
Multimodal model training. Small multimodal models (e.g., CLIP [61] and LiT [62]) have been widely studied in recent years. Many system optimizations have been proposed to train such multimodal models efficiently. DistMM [63] tackles model heterogeneity by introducing modality-aware placement and partitioning to evenly distribute workload. GraphPipe [64] presents graph pipeline parallelism, addressing graph dependencies in multimodal models to minimize pipeline bubbles. Yet, these advancements primarily enhance small multimodal models trained on tens of GPUs. They fall short for scaling up to meet the demands of integrating multimodal models with LLMs which necessitate training across thousands of GPUs. This gap underpins the motivation behind MMScale, designed to meet the unique challenges of multimodal large language model training.
Multimodal LLM serving. LLM serving has been widely studied in recent years. Orca [65], FastServe [66], and VTC [67] propose iteration-level scheduling algorithms to improve the serving quality. DistServe [68] and Splitwise [69] propose disaggregated serving for prefill and decoding to improve the serving throughput. vLLM [70], RAGCache [71], and SGLang [72] propose prefix caching to reuse the KV tensors. However, the serving of multimodal LLMs remains under-explored. MMScale’s core insights (e.g., disaggregated modules) are applicable to multimodal LLMs serving. Our future work will delve into the specific challenges posed by multimodal LLMs serving systems.
10 Conclusion
We present MMScale, an efficient and adaptive framework to reform the training of multimodal LLMs. We identify the key challenges in training multimodal LLMs, i.e., model heterogeneity and data heterogeneity. MMScale introduces novel adaptive resource allocation to eliminate model heterogeneity and data-aware reordering to mitigate data heterogeneity. We evaluate MMScale on production cluster with thousands of GPUs and show that it achieves 54.7% MFU and outperforms Megatron-LM by up to 2.2 on throughput.

### v2

@@
-Addressing Model and Data Heterogeneity in Multimodal Large Language Model Training
+DistTrain: Addressing Model and Data Heterogeneity with Disaggregated Training for Multimodal Large Language Models
 Abstract
@@
 Yet, training multimodal LLMs suffers from low efficiency and scalability, due to the inherent model heterogeneity and data heterogeneity across different modalities.
-We present MMScale, an efficient and adaptive framework to reform the training of multimodal large language models on large-scale clusters.
-MMScale exploits the system characteristics of multimodal LLM training to achieve high efficiency and scalability.
-The core of MMScale is the adaptive resource allocation and data-aware reordering techniques to eliminate the model and data heterogeneity respectively.
-We also tailor system optimizations for multimodal LLM training to offload certain operations from the GPU training.
-We evaluate MMScale across different sizes of multimodal LLMs on a large-scale production cluster with thousands of GPUs.
-The experimental results show that MMScale achieves 54.7% Model FLOPs Utilization (MFU) when training a 72B multimodal LLM on 1172 GPUs and outperforms Megatron-LM by up to 2.2 on throughput.
-The ablation study shows the main techniques of MMScale are both effective and lightweight.
+We present DistTrain, an efficient and adaptive framework to reform the training of multimodal large language models on large-scale clusters.
+The core of DistTrain is the disaggregated training technique that exploits the characteristics of multimodal LLM training to achieve high efficiency and scalability.
+Specifically, it leverages disaggregated model orchestration and disaggregated data reordering to address model and data heterogeneity respectively.
+We also tailor system optimization for multimodal LLM training to overlap GPU communication and computation.
+We evaluate DistTrain across different sizes of multimodal LLMs on a large-scale production cluster with thousands of GPUs.
+The experimental results show that DistTrain achieves 54.7% Model FLOPs Utilization (MFU) when training a 72B multimodal LLM on 1172 GPUs and outperforms Megatron-LM by up to 2.2 on throughput.
+The ablation study shows the main techniques of DistTrain are both effective and lightweight.
 1 Introduction
@@
 Subsequently, the modality generator translates this processed information back into coherent outputs tailored to each modality.
-Megatron-LM [17] is the state-of-the-art framework for training LLMs, which comprise a stack of homogeneous transformer layers.
-It can be extended to train multimodal LLMs by treating the multimodal modules as additional layers within the LLM.
-However, training multimodal LLMs with Megatron-LM poses two substantial challenges: model heterogeneity and data heterogeneity.
+Existing LLM training framework, e.g., Megatron-LM [17], can be extended to train multimodal LLMs by treating the multimodal modules as additional layers within the LLM.
+However, training multimodal LLMs poses two substantial challenges: model heterogeneity and data heterogeneity.
 The fundamental issue of model heterogeneity stems from the need to process diverse modalities with different modules that vary dramatically in size and operator complexity.
@@
 These challenges collectively limit the efficiency and scalability of multimodal LLM training, resulting in MFU as low as ~20% in production-level training (§8.1).
-To this end, we present MMScale, an efficient and adaptive framework to reform the training of multimodal LLMs.
-MMScale achieves state-of-the-art MFU which is close to unimodal (i.e., text-only) LLM training and effectively scales to large clusters with thousands of GPUs.
-The core of MMScale is adaptive resource allocation and data-aware reordering.
-These two techniques leverage the distinctive system characteristics of multimodal LLM training to eliminate model heterogeneity and data heterogeneity respectively.
+To this end, we present DistTrain, an efficient and adaptive framework to reform the training of multimodal LLMs.
+DistTrain achieves state-of-the-art MFU which is close to unimodal LLM training and effectively scales to large clusters with thousands of GPUs.
+The core principle of DistTrain is disaggregated training, including GPU training disaggregation and CPU preprocessing disaggregation, which facilitate to address model and data heterogeneity respectively.
 For model heterogeneity, we meticulously analyze the pipeline bubbles stemming from model heterogeneity and identify their root causes (§2.2).
-Based on this analysis, we propose disaggregated training, which tailors the resource allocation and parallelism strategies across the modality encoder, LLM backbone, and modality generator.
-We also develop a novel and lightweight algorithm to navigate this complicated design space to choose the optimal resource and parallelism configurations.
-This innovative approach minimizes the pipeline bubbles and achieves optimal training efficiency.
-For data heterogeneity, we categorize the training stragglers into inter-microbatch and intra-microbatch stragglers (§2.3) and introduce tailored algorithms that strategically reorder training data while maintaining synchronous training semantics.
+The GPU training disaggregation of modality encoder, LLM backbone, and modality generator enables adaptive orchestration across the three modules in multimodal LLM.
+Building on this, we propose disaggregated model orchestration that navigates the complicated design space to choose the optimal resource and parallelism configurations.
+This innovative approach minimizes the pipeline bubbles caused by model heterogeneity and achieves optimal training efficiency.
+For data heterogeneity, we categorize the training stragglers into inter-microbatch and intra-microbatch stragglers (§2.3).
+The CPU preprocessing disaggregation allows efficient data preprocessing (e.g., data decompression and reordering) with negligible runtime overhead.
+From this foundation, we incorporate disaggregated data reordering into the preprocessing to strategically reorder training data without additional overhead.
 The inter-microbatch reordering algorithm reorders the data samples to evenly distribute the load across different data parallelism groups.
 The intra-microbatch reordering algorithm reorders the microbatches tailored for (interleaved) 1F1B pipeline scheme to minimize the pipeline bubbles.
-This two-level approach effectively mitigates training stragglers caused by data heterogeneity.
-In addition, we customize system optimizations for multimodal LLM training to offload certain operations from the critical path (i.e., GPU computation).
-We implement an asynchronous data loader to pre-process the training data with negligible runtime overhead.
+This two-level approach effectively mitigates data heterogeneity.
+In addition, we customize system optimization for multimodal LLM training.
 We implement an in-house collective communication library, StepCCL, to hide the communication overhead within the computation.
 In summary, we make the following contributions.
-We present MMScale, an efficient and adaptive framework to reform the training of multimodal LLMs by harnessing the multimodal heterogeneity.
+We present DistTrain, an efficient and adaptive framework to reform the training of multimodal LLMs by addressing the multimodal heterogeneity.
 It delivers state-of-the-art MFU on large-scale clusters with thousands of GPUs.
 We identify and discuss the primary challenges associated with multimodal LLM training, which are summarized as model heterogeneity and data heterogeneity.
-We implement disaggregated training across the three modules in multimodal LLMs.
-We propose adaptive resource allocation and data-aware reordering to effectively address model and data heterogeneity, respectively.
-We implement MMScale and conduct experiments on our production cluster with thousands of GPUs.
-The experimental results show that MMScale achieves 54.7% MFU when training a 72B multimodal LLM on 1172 GPUs and outperforms Megatron-LM by up to 2.2 on throughput.
+We propose disaggregated training for multimodal LLM.
+It leverages disaggregated model orchestration and disaggregated data reordering to effectively address model and data heterogeneity, respectively.
+We implement DistTrain and conduct experiments on our production cluster with thousands of GPUs.
+The experimental results show that DistTrain achieves 54.7% MFU when training a 72B multimodal LLM on 1172 GPUs and outperforms Megatron-LM by up to 2.2 on throughput.
 2 Motivation
@@
 The output data of the LLM backbone is subsequently refined by output projection layers, which tailor the information for each modality.
-Finally, the modality generator (e.g., Diffusion [27] for images and AudioLDM [28] for audios) transforms the LLM-processed information back into the respective modal outputs to generate coherent and contextually relevant results.
+Finally, the modality generator (e.g., Diffusion [27] for images and AudioLDM [28] for audios) transforms the LLM-processed information back into the respective modal outputs for generation.
 Multimodal LLM training necessitates training all three modules simultaneously.
@@
 The projector is co-located with the modality encoder and generator and is replicated across the GPUs in the TP group.
-However, this training framework introduces significant computation imbalance stemming from model heterogeneity and data heterogeneity due to its rigid resource allocation method (i.e., the multimodal modules share the same DP and TP strategies with LLM backbone).
+However, this training framework introduces significant computation imbalance stemming from model heterogeneity and data heterogeneity due to its rigid model orchestration method.
+(i.e., the multimodal modules share the same DP and TP strategies with LLM backbone).
 It only achieves ~20% MFU (§8.1) which is significantly lower than the MFU of ~50% observed in training unimodal (text-only) LLMs [29].
@@
 The straggler (i.e., the microbatch ) significantly delays the training process of the subsequent PP stages, leading to a large pipeline bubble.
-3 MMScale Overview
-We present MMScale, an efficient and adaptive framework to reform the training of multimodal LLMs by harnessing the multimodal heterogeneity.
-MMScale exploits the characteristics of parallelism training of multimodal LLMs to achieve high efficiency and scalability.
-MMScale eliminates the model heterogeneity by adaptive resource allocation (§4) and harnesses the data heterogeneity by data-aware reordering (§5).
-In addition, MMScale adopts some system optimizations tailored for multimodal LLM training (§6).
-Here we provide a brief overview of MMScale as Figure 8 shows.
-MMScale orchestrator.
-Before training, MMScale employs a training orchestrator to determine the resource allocation and parallelism strategy for each module in multimodal LLMs.
+3 DistTrain Overview
+We present DistTrain, an efficient and adaptive framework to reform the training of multimodal LLMs by addressing the multimodal heterogeneity.
+DistTrain proposes disaggregated training for multimodal LLM to achieve high efficiency and scalability.
+DistTrain eliminates the model heterogeneity by disaggregated model orchestration (§5) and harnesses the data heterogeneity by disaggregated data reordering (§6).
+In addition, DistTrain adopts some system optimizations tailored for multimodal LLM training (§7).
+Here we provide a brief overview of DistTrain as Figure 8 shows.
+DistTrain manager.
+Before training, DistTrain employs a training manager to determine the resource allocation and parallelism strategy for each module in multimodal LLMs.
 This scheduler first gathers the model architecture and training configuration (e.g., global batch size) from the user and randomly samples a subset of training data to analyze the data distribution.
 Utilizing this information, it runs a series of benchmarking training trials and constructs a performance profiler with linear interpolation to estimate each module’s computation and communication time.
-Based on the profiling results, the training orchestrator decides the optimal resource allocation and parallelism strategy with adaptive resource allocation for one specific training task, as detailed in §4.
-MMScale initializer.
-Then, MMScale initializes the parallelism units for the modality encoder, LLM backbone, and modality generator, respectively.
-MMScale allocates different numbers of GPUs to each unit.
-Each unit then establishes the communication group across different nodes and GPUs.
+Based on the profiling results, the training manager decides the optimal resource allocation and parallelism strategy with disaggregated model orchestration for one specific training task, as detailed in §5.
+DistTrain initializer.
+DistTrain then initializes the GPU training disaggregation for modality encoder, LLM backbone, and modality generator, respectively.
+DistTrain allocates different numbers of GPUs to each parallelism unit.
+Each unit then establishes the specific communication group.
 The unit loads the model checkpoint from distributed file system and shards the model parameters and optimization states.
-Finally, MMScale conducts several communication trials to warm up the system and test connectivity.
-MMScale runtime.
-After initialization, MMScale begins the main training process.
-In each training iteration, it first retrieves a global batch of training samples from the distributed file system for pre-processing.
-Then, MMScale performs data-aware reordering to reorder the training samples within one global batch without breaking the synchronous training semantics [31].
-This module effectively eliminates both inter-microbatch and intra-microbatch data heterogeneity, as detailed in §5.
+Finally, DistTrain conducts several communication trials to warm up the system and test connectivity.
+DistTrain runtime.
+At runtime, the dedicated CPU nodes (i.e., CPU preprocessing disaggregation) retrieve training samples from the distributed file system for preprocessing.
+It performs disaggregated data reordering to reorder the training samples within one global batch without breaking the synchronous training semantics [31].
+This reordering effectively eliminates both inter-microbatch and intra-microbatch data heterogeneity, as detailed in §6.
+In each iteration, the main training process receives the preprocessed data asynchronously from the CPU nodes.
 The data then undergoes sequentially through the modality encoder, LLM backbone, and modality generator in the training pipeline.
 Finally, it synchronizes the gradients and model parameters through the all-gather operation, employing the ZERO-1 optimization [32] and mixed precision training [33].
-Additionally, MMScale adopts a dedicated process to periodically and asynchronously save the model checkpoint to the distributed file system for fault tolerance.
-4 Addressing Model Heterogeneity
-Before diving into the details, we first introduce the disaggregation of the training across the three modules (i.e., modality encoder, LLM backbone, and modality generator) in multimodal LLMs.
-This disaggregation facilitates adaptive GPU resource allocation among the different modules based on the training workload.
-Subsequently, we formulate the problem of adaptive resource allocation to minimize the training time per iteration.
-Finally, we present the detailed algorithm to optimally address the problem caused by model heterogeneity.
-4.1 Disaggregated Training
-Disaggregating the training among the three modules is the foundation of adaptive resource allocation in MMScale.
-Figure 9 demonstrates the training topology of disaggregated training in MMScale.
-Different from the rigid resource allocation in Megatron-LM (i.e., Figure 2), MMScale is able to adaptively adjust the resource allocation.
-For instance, MMScale allocates 4 GPUs (DP=2 and TP=2) to the modality encoder, 12 GPUs (DP=3 and TP=4) to the LLM backbone per PP stage, and 4 GPUs (DP=1 and TP=4) to the modality generator.
-Additionally, the projector layers are colocated with either the modality encoder or generator, with their number of replicas adapting as needed.
-We implement the disaggregated training through a dedicated module, i.e., parallelism unit.
+Additionally, DistTrain adopts a dedicated process to periodically and asynchronously save the model checkpoint to the distributed file system for fault tolerance.
+4 Disaggregated Training
+To address the model and data heterogeneity in multimodal LLM training, we first introduce DistTrain’s core principle: disaggregated training.
+It includes GPU training disaggregation and CPU preprocessing disaggregation.
+GPU training disaggregation provides opportunities for adaptive model orchestration across the three modules to address the model heterogeneity.
+CPU preprocessing disaggregation facilitates data preprocessing to address data heterogeneity with negligible runtime overhead.
+4.1 GPU Training Disaggregation
+Figure 9 demonstrates the training topology of disaggregated GPU training in DistTrain.
+Different from the rigid model orchestration in Megatron-LM (i.e., Figure 2), DistTrain is able to adaptively adjust the resource allocation and parallelism strategy.
+For instance, DistTrain allocates 4 GPUs (DP=2 and TP=2) to the modality encoder, 12 GPUs (DP=3 and TP=4) to the LLM backbone per PP stage, and 4 GPUs (DP=1 and TP=4) to the modality generator.
+Additionally, the projector layers are co-located with either the modality encoder or generator, with their number of replicas adapting as needed.
+We implement GPU training disaggregation through a dedicated module, i.e., parallelism unit.
 Parallelism unit.
 At training initialization, we need to establish the communication group according to the resource allocation and parallelism strategy.
-MMScale introduces a module, parallelism unit, composed of one or more PP stages.
+DistTrain introduces a module, parallelism unit, composed of one or more PP stages.
 Each unit can adopt its own DP and TP strategies and form a specific communication group.
 Inter-unit connections are facilitated by a communication broker, which bridges PP communication across parallelism units.
-Users are only required to specify the DP and TP configurations for each parallelism unit, and MMScale automatically sets up the communication group and communication broker.
-MMScale treats the modality encoder, LLM backbone, and modality generator as three individual parallelism units.
+Users are only required to specify the DP and TP configurations for each parallelism unit, and DistTrain automatically sets up the communication group and communication broker.
+DistTrain treats the modality encoder, LLM backbone, and modality generator as three individual parallelism units.
 The detailed implementation of parallelism unit is discussed in §7.
+4.2 CPU Preprocessing Disaggregation
+When training multimodal LLMs, training samples often combine lightweight text with heavyweight multimodal data.
+The latter significantly increases data preprocessing time.
+For example, a typical training sample could include a 256-word text sequence and ten 10241024 RGB images.
+The text is just kilobytes, whereas the images are total of 120 megabytes.
+Preprocessing (e.g., decompression, resizing, and reordering) such samples can take several seconds and interfere with the co-located training process.
+DistTrain disaggregates the CPU data preprocessing from the GPU training process with a producer-consumer model.
+The producer, operating on dedicated CPU nodes, fetches data from the distributed file system and preprocesses training data asynchronously with the GPU training process.
+The consumer, i.e., the main training process, receives this preprocessed data for training.
+The producer and consumer communicate through RPC calls, and use RDMA network for lower latency if available.
+This disaggregation guarantees that the CPU preprocessing does not interfere with the GPU training process and achieves negligible data preprocessing overhead.
+5 Addressing Model Heterogeneity
+Disaggregated training enables disaggregated model orchestration among the different modules based on the training workload.
+We first formulate the problem of disaggregated model orchestration to minimize the training time per iteration.
+Then, we present the detailed algorithm to optimally address the problem caused by model heterogeneity.
 5.1 Problem Formulation
-With disaggregated training, we are able to adaptively allocate GPU resources across the three modules.
+With disaggregated training, we are able to adaptively orchestrate the three modules.
 The problem now lies in determining the optimal resource allocation and parallelism strategy to minimize the training time per iteration.
-Exhaust search is infeasible due to the large search space of the parallelism strategy and cluster size.
-One strawman solution is to allocate the GPU resources proportional to the model flops of each module.
+Exhaust search is infeasible due to the large search space, particularly in large cluster.
+One strawman solution is to allocate the resources proportional to the model flops of each module.
 However, this method falls short as it overlooks complex patterns in parallelism training.
-For instance, different DP and TP sizes in modality encoder could lead to different computation times while the model flops are the same.
-Before diving into the adaptive resource allocation algorithm, we first formulate the optimization problem.
+Before diving into disaggregated model orchestration, we first formulate the optimization problem.
 LLM backbone.
@@
 Modality encoder and generator.
-In MMScale, the modality encoder is regarded as a parallelism unit with PP size .
+In DistTrain, the modality encoder is regarded as a parallelism unit with PP size .
 Let the TP size be and the DP size be .
@@
 The forward time of one PP stage for one microbatch in the modality encoder is .
-Similarly, the forward time of one PP stage in the modality generator is , where is the number of GPUs allocated to modality encoder.
+Similarly, the forward time of one PP stage in the modality generator is , where is the number of GPUs allocated to modality generator.
 Objective function.
@@
 We will retrofit the formulation to adapt to VPP later.
-As for the communication time, it consists of the time of P2P communication in PP and collective communication in DP.
-It is modeled as the communication volume divided by the communication bandwidth.
-If the sizes of TP and DP are fixed, the communication time remains constant.
+TP communication is incorporated into the functions , and , which are calibrated through interpolation from actual trials.
+The communication time of DP and PP is modeled as the communication volume divided by the bandwidth.
+If the DP sizes are fixed, the DP communication time remains constant.
+For PP, the communication time equals the PP size multiplied by the communication time of a single layer’s output tensor.
 Constraints.
@@
 In 1F1B, the first PP stage requires storage for microbatches of activation states.
-We eschew using GPipe in MMScale since GPipe consumes more memory.
+We eschew using GPipe in DistTrain since GPipe consumes more memory.
 The memory constraint ensures the sum of the four memory parts on one GPU doesn’t exceed GPU capacity.
 As for modality encoder and generator, the formulation is similar.
-4.3 Adaptive Resource Allocation
+5.2 Disaggregated Model Orchestration
 The optimization problem is non-convex, with , and the DP, TP sizes as positive variables.
@@
 Designing an efficient algorithm that quickly identifies the optimal resource allocation and parallelism strategy is a significant challenge.
-Convex optimization on resource allocation.
-Our key insight is to decompose the non-convex optimization problem into a series of simplified convex problems with variables (i.e., the resource allocation).
+Convex optimization.
+Our key insight is to decompose the non-convex optimization problem into a series of simplified convex problems with variables (i.e., resource allocation).
 We confine the TP size to on an NVIDIA GPU node with 8 GPUs and adjust the DP size as a factor of to balance the computation across DP groups.
 The PP size of the LLM backbone is calculated as .
-The set of possible parallelism strategies is a manageable and finite set, i.e., the Cartesian product of the TP and DP size.
+The set of possible parallelism strategies is a manageable and finite set, i.e., the Cartesian product of TP and DP size.
 This allows us to enumerate all feasible TP and DP sizes and transform the original optimization problem into a set of simplified problems.
@@
 As a result, the simplified optimization problem is convex and can be efficiently solved to optimality by existing solvers [34, 35].
-The algorithm, detailed in Algorithm 1, efficiently finds the optimal resource allocation and parallelism strategy to minimize iteration time.
+The algorithm, detailed in Algorithm 1, efficiently finds the optimal resource allocation and parallelism strategy.
 Virtual pipeline parallelism (i.e., interleaved 1F1B) reduces the warm-up time by dividing model into finer-grained virtual PP (VPP) stages.
@@
 6 Addressing Data Heterogeneity
-The data heterogeneity causes intra-microbatch and inter-microbatch stragglers.
-Addressing both straggler simultaneously renders the problem intractable due to its high complexity (e.g., the order of the training samples results in factorial-level complexity) and the necessity of runtime execution.
-To tackle this challenge, we decompose the overarching problem into two tractable subproblems.
+Disaggregated training enables preprocessing the training data with negligible runtime overhead.
+Based on this, we introduce disaggregated data reordering, seamlessly integrated into data preprocessing, to address data heterogeneity without additional overhead.
 Initially, we present intra-microbatch reordering to eliminate stragglers across DP groups.
 We then introduce inter-microbatch reordering to minimize pipeline bubbles caused by the straggler microbatches.
-The combination of the two reordering, i.e., data-aware reordering, effectively addresses the data heterogeneity.
+The combination of the two reordering, i.e., disaggregated data reordering, effectively addresses the data heterogeneity.
+We emphasize that these two reordering algorithms are integrated into the disaggregated data preprocessing.
+This ensures that the complex reordering does not interfere with the GPU training process.
 6.1 Intra-microbatch Reordering
 Insight.
-To address the first subproblem of intra-microbatch stragglers, we identify the straggler by pinpointing the DP group with the largest training samples.
+To address the first subproblem, intra-microbatch stragglers, we identify the straggler by pinpointing the DP group with the largest training samples.
 As illustrated in Figure 6, the first DP group becomes a straggler as it contains the two largest training samples.
@@
 There is no known polynomial-time algorithm capable of determining the optimal solution.
-Given the variability of runtime training samples, any algorithm employed must run before each iteration with negligible runtime overhead.
+Given the substantial batch size in production training, the algorithm employed must be lightweight and efficient.
 Consequently, we adopt the greedy number partitioning algorithm, which guarantees an approximation ratio of [37].
@@
 Then, it loops over the training samples and assigns the sample to the DP group with the current lowest computational load (line 6-8).
-It then returns the reordered training samples (line 9-11).
-The algorithm is lightweight and maintains a time complexity of .
+It then returns the reordered samples (line 9-11).
+The algorithm has a time complexity of .
 6.2 Inter-microbatch Reordering
@@
 Straggler microbatches in either the encoder or generator prolong these intervals or increase the unfilled area (i.e., bubble).
-Designing an efficient runtime algorithm to minimize pipeline bubbles presents a significant challenge.
 Insights.
 We leverage two insights to solve this problem.
-The first insight involves minimizing the volume of intervals that cannot be filled.
+The first insight involves minimizing the volume of intervals that are not filled.
 As shown in Figure 12, the last intervals (i.e., to ) remain unfilled.
@@
 We observe a positive correlation between the volume of and the size of the microbatch.
-For instance, is significantly larger than and .
+The size refers to the computation time of the microbatch in modality encoder and generator.
+For instance, is significantly larger than and since the microbatch is the largest.
 By strategically reordering the training samples to position the smallest microbatches at the end, we are able to reduce these unfilled intervals (i.e., pipeline bubbles).
-The second insight involves minimizing the unfilled area of front intervals.
-As shown in Figure 12, the first interval is filled with by the to microbatches’ forward passes.
-The volume of the first interval (i.e., ) is determined by the aggregate forward and backward times of the first microbatch in LLM backbone and modality generator.
-Such characteristic allows the accommodation of large microbatches early in the sequence to fill the first interval.
-For subsequent intervals, is filled by the microbatch’s forward pass.
-By evaluating the volume of , we place the microbatch, whose forward time most closely matches this volume, at the position, thereby reducing the unfilled area (i.e., pipeline bubbles).
+The second insight involves minimizing the unfilled area of left intervals.
+The left intervals (i.e., to ) are filled with the forward pass.
+As shown in Figure 12, the first interval is filled with by the to forward passes.
+For subsequent intervals, is filled by the forward pass.
+By evaluating the volume of , we place the microbatches, whose forward time most closely matches this volume, at the corresponding position, to minimize the unfilled area (i.e., pipeline bubbles).
 Inter-microbatch Reordering.
@@
 This dynamic programming algorithm exhibits a complexity of per function invocation.
-The algorithm is lightweight and efficient with a time complexity of .
+The algorithm has a time complexity of .
 Virtual pipeline parallelism (i.e., interleaved 1F1B) also follows the one forward and one backward pipeline scheme to reduce the memory footprint.
@@
 7 System Implementation and Optimization
-We implement MMScale with 6.3K lines of code in Python and C++, and integrate it with Megatron-LM [17], a state-of-the-art training framework for large language models.
-MMScale is able to support multimodal LLM training with a variety of modalities on a large-scale GPU cluster (e.g., thousands of GPUs used in §7).
-MMScale leverages a distributed file system to store the training data and model checkpoints.
+We implement DistTrain with 6.3K lines of code in Python and C++, and integrate it with Megatron-LM [17], a state-of-the-art training framework for large language models.
+DistTrain is able to support multimodal LLM training with a variety of modalities on a large-scale GPU cluster (e.g., thousands of GPUs used in §8).
+DistTrain leverages a distributed file system to store the training data and model checkpoints.
 It handles failures by automatically recovering the training process from the latest checkpoint.
-MMScale orchestrator.
-MMScale’s training orchestrator, implemented as a Python script on a dedicated CPU node, formulates the adaptive resource allocation problem using Disciplined Convex Programming [38].
+DistTrain manager.
+DistTrain’s training manager, implemented as a Python script on a dedicated CPU node, formulates the disaggregated model orchestration problem using Disciplined Convex Programming [38].
 It employs the CVX solver [39] to efficiently solve this problem within milliseconds.
-The orchestrator then records the optimal resource allocation and parallelism strategy to a configuration file, which the Kubernetes controller uses to launch the training task.
-MMScale initializer.
-MMScale incorporates parallelism unit that manages specific parallelism strategies for each module in multimodal LLMs.
-Parallelism unit is initialized using PyTorch Distributed [40] library with NCCL as the default communication backend, except for the TP communication where we use StepCCL instead.
-MMScale shards the model parameters in accordance with the established parallelism strategy.
+The manager then records the optimal resource allocation and parallelism strategy to a configuration file, which the Kubernetes controller uses to launch the training task.
+DistTrain initializer.
+DistTrain incorporates parallelism unit that manages specific parallelism strategies for each module in multimodal LLMs.
+Parallelism unit is initialized using PyTorch Distributed [40] library with NCCL as the communication backend, except for the TP communication where we use StepCCL instead.
+DistTrain shards the model parameters in accordance with the established parallelism strategy.
 Parallelism unit.
-As we discussed in §4.1, the disaggregated training is implemented through a specific module: parallelism unit.
-When initializing the distributed training, MMScale first establishes the communication groups within one parallelism unit.
+As we discussed in §4.1, the GPU training disaggregation is implemented through a specific module: parallelism unit.
+When initializing the distributed training, DistTrain first establishes the communication groups within one parallelism unit.
 Each GPU process has a global rank and a local rank within the unit, which facilitates the distributed initialization.
@@
 To alleviate this, we implement asynchronous send operations that eliminate these superfluous dependencies and redesign the communication topology to prevent potential deadlocks.
-Asynchronous data loader.
-When training multimodal LLMs, training samples often combine lightweight text with heavyweight multimodal data.
-The latter significantly increases data loading time.
-For example, a typical training sample could include a 256-word text sequence and ten 10241024 RGB images.
-The text is just kilobytes, whereas the images are total of 120 megabytes.
-Loading multiple such samples can take several seconds.
-MMScale’s data loader employs a producer-consumer model to enhance efficiency.
-The producer, operating on dedicated CPU nodes, fetches data from the distributed file system and pre-processes training data (e.g., image cropping) asynchronously with the GPU training process.
-The consumer, integrated within the main training process, receives this pre-processed data and reorders it to eliminate the data heterogeneity.
-The producer and consumer communicate through RPC calls, and use RDMA network for lower latency if available.
 Mitigating TP overhead with StepCCL.
@@
 8 Evaluation
-In this section, we first use large-scale experiments to demonstrate the overall performance improvements of MMScale over existing multimodal LLM training solution.
-Next, we use microbenchmarks to deep dive into MMScale and show the effectiveness of each component of MMScale.
-Finally, we conduct a case study to evaluate MMScale under different frozen settings and evaluate the system optimizations.
+In this section, we first use large-scale experiments to demonstrate the overall performance improvements of DistTrain over Megatron-LM.
+Next, we use microbenchmarks to deep dive into DistTrain and show the effectiveness of each component in DistTrain.
+Finally, we provide a case study to further elucidate DistTrain’s capabilities.
 | Models | # of | Hidden | FFN Hidden | # of | # of |
@@
 As for the modality, we focus on images and texts.
-MMScale is also compatible with other modalities.
+DistTrain is also compatible with other modalities.
 For modality encoder and modality generator, we use ViT-Huge [45] (0.63B) and Stable-Diffusion 2.1 [46] (1B) respectively.
@@
 Metrics.
-We use the Model FLOPs Utilization (MFU) as the primary metric to evaluate MMScale.
+We use the Model FLOPs Utilization (MFU) as the primary metric to evaluate DistTrain.
 MFU measures the percentage of GPU FLOPs that are effectively utilized during model training.
-We also use the training throughput (TPT.) to evaluate the training speed of MMScale.
-Since MMScale and Megatron-LM may utilize different numbers of GPUs due to varying resource allocation strategies, we also indicate the number of GPUs used in the throughput charts.
+We also use the training throughput (TPT.) to evaluate the training speed of DistTrain.
+Since DistTrain and Megatron-LM may utilize different numbers of GPUs due to varying model orchestration strategies, we also indicate the number of GPUs used in the throughput charts.
 8.1 Overall Performance
 Setup.
-We first compare the overall performance of MMScale against Megatron-LM on a large-scale GPU cluster (up to 1296 GPUs).
+We first compare the overall performance of DistTrain against Megatron-LM on a large-scale GPU cluster (up to 1296 GPUs).
 We retrofit Megatron-LM to support multimodal LLM training by integrating modality encoder and generator into the training pipeline.
-Megatron-LM employs rigid resource allocation as described in §2.1.
-In Megatron-LM, we set the PP size of the LLM backbone to 1, 2, and 10 for Llama3-7B, Llama3-13B, and Llama3-70B, respectively.
+Megatron-LM employs rigid model orchestration as described in §2.1.
+In Megatron-LM, we set the PP size of the LLM backbone to 1, 2, and 10 for Llama3-7B, Llama3-13B, and Llama3-70B.
+PP size is set to 1 for modality encoder and generator.
 TP size is set to 8.
-As for MMScale, the parallelism strategy is determined by adaptive resource allocation.
+As for DistTrain, the parallelism strategy is determined by disaggregated model orchestration.
 In our experiments, one GPU is able to facilitate training ViT and Stable-Diffusion.
-For both Megatron-LM and MMScale, we replicate the encoder and generator across the GPUs within the TP group to process different image subsequences, whereas TP itself is not used.
+We replicate the encoder and generator across the GPUs within the TP group to process different images, whereas TP itself is not used.
 We set global batch size to 1920.
 The experimental results are shown in Figure 15 and Figure 16.
-Figure 15 shows the MFU of MMScale and Megatron-LM under different multimodal LLMs.
+Figure 15 shows the MFU.
 Figure 16 shows the training throughput and marks the number of GPUs used in each experiment.
+The different GPU numbers are due to the varying model orchestration strategies of DistTrain and Megatron-LM.
 We summarize the experiment as follows.
-As shown in Figure 15, MMScale achieves 51.8%-54.7% MFU under large-scale multimodal LLM training scenario.
-This performance closely approximates that of state-of-the-art unimodal (i.e., text) LLM training [29], which demonstrates the effectiveness of MMScale in eliminating the model and data heterogeneity in multimodal LLM training.
-MMScale significantly outperforms Megatron-LM, delivering 1.7-2.8 the MFU and 1.7-2.2 the training throughput when training MLLM-9B and MLLM-15B with a similar number of GPUs.
-These performance gains largely stem from MMScale’s adaptive resource allocation strategy.
-Megatron-LM’s rigid strategy often leads to GPU underutilization in multimodal modules, since it assigns too many GPUs to the modality encoder and generator.
-In contrast, MMScale adaptively adjusts GPU allocations based on specific model and data demands.
-Additionally, MMScale’s data-aware reordering technique further boosts efficiency.
-In MLLM-72B training scenario, MMScale also outperforms Megatron-LM by 1.2 on MFU and 1.3 on training throughput with a similar number of GPUs.
+As shown in Figure 15, DistTrain achieves 51.8%-54.7% MFU in large-scale multimodal LLM training.
+This performance closely approximates that of state-of-the-art unimodal (i.e., text) LLM training [29], which demonstrates the effectiveness of DistTrain in addressing the model and data heterogeneity in multimodal LLM training.
+DistTrain significantly outperforms Megatron-LM, delivering 1.7-2.8 the MFU and 1.7-2.2 the training throughput when training MLLM-9B and MLLM-15B with a similar number of GPUs.
+These performance gains largely stem from DistTrain’s disaggregated model orchestration.
+Megatron-LM’s rigid strategy often leads to GPU underutilization, since it assigns too many GPUs to the modality encoder and generator.
+In contrast, DistTrain adaptively adjusts model orchestration based on specific model and data demands.
+Additionally, DistTrain’s disaggregated data reordering technique further boosts efficiency.
+In MLLM-72B training scenario, DistTrain also outperforms Megatron-LM by 1.2 on MFU and 1.3 on training throughput with a similar number of GPUs.
 The high image resolution prolongs the execution time of the multimodal module, which introduces pipeline bubbles in LLM backbone.
-MMScale addresses this by allocating additional GPUs to these modules to balance the pipeline.
-MMScale’s data-aware reordering strategy continues to diminish data heterogeneity, thereby increasing overall training efficiency.
-7.2 Deep Dive into MMScale
-In this subsection, we perform microbenchmarks to evaluate the effectiveness of each MMScale’s component.
+DistTrain addresses this by allocating additional GPUs to these modules to balance the pipeline.
+The disaggregated data reordering strategy continues to diminish data heterogeneity, thereby increasing training efficiency.
+8.2 Deep Dive into DistTrain
+In this subsection, we perform microbenchmarks to evaluate the effectiveness of each DistTrain’s component.
 Our experiments utilize up to 98 NVIDIA A800 GPUs.
 We set the global batch size to 128, 64, and 40 for MLLM-9B, MLLM-15B, and MLLM-72B, respectively.
-7.2.1 Adaptive Resource Allocation
-We measure the MFU and training throughput of MMScale and other resource allocation strategies, including a comparison with Megatron-LM and a ratio-based approach.
+8.2.1 Disaggregated Model Orchestration
+We measure the MFU and training throughput of DistTrain and other model orchestration strategies, including a comparison with Megatron-LM and a ratio-based approach.
 The ratio-based strategy allocates GPUs according to the computational demands (flops) of each module.
@@
 The experimental results are shown in Figure 17.
-MMScale consistently outperforms the baseline strategies, achieving 1.3%-2.7% higher MFU and 1.4%-2.7% higher training throughput.
-Although the ratio-based strategy outperforms Megatron-LM’s rigid allocation, it still lags behind MMScale since it neglects the intricate performance model (§4.2) of multimodal LLM training.
-MMScale employs an adaptive resource allocation strategy that optimally balances computational loads across the three modules and achieves high resource utilization.
-We also evaluate the running time of MMScale’s resource allocation algorithm under different training settings, as detailed in Table 2.
-The algorithm completes the allocation process in under one second.
+DistTrain consistently outperforms the baseline strategies, achieving 1.3-2.7 higher MFU and 1.4-2.7 higher training throughput.
+Although the ratio-based strategy outperforms Megatron-LM’s rigid strategy, it still lags behind DistTrain since it neglects the intricate performance model (§5.1) of multimodal LLM training.
+DistTrain’s disaggregated model orchestration optimally balances computational loads across the three modules and achieves high resource utilization.
+(a) MFU. (b) Throughput.
+We also evaluate the running time of DistTrain’s disaggregated model orchestration under different training settings, as detailed in Table 2.
+The algorithm completes in under one second.
 The overhead is negligible compared to the days or even weeks required for overall training.
-7.2.2 Data-aware Reordering
-We evaluate the effectiveness of MMScale’s data-aware reordering by comparing it against the random order set by PyTorch’s DataLoader, while keeping other components the same.
+8.2.2 Disaggregated Data Reordering
+We evaluate the effectiveness of DistTrain’s disaggregated data reordering by comparing it against the random order, while keeping other components the same.
 The effectiveness is gauged through metrics such as MFU and training throughput (TPT.).
-We use the optimal resource allocation and parallelism strategy decided by MMScale’s adaptive resource allocation algorithm.
-Given that the GPU allocation policy remains unchanged, the number of GPUs is not specified in this experiment.
+We use the optimal resource allocation and parallelism strategy decided by DistTrain’s disaggregated model orchestration.
+Given that the model orchestration strategy remains unchanged, the number of GPUs is not shown.
 The experimental settings are the same as those in §8.2.1.
 The results are shown in Figure 18.
-MMScale consistently outperforms the baseline, achieving 1.03%-1.11% higher MFU and training throughput.
+DistTrain consistently outperforms the baseline, achieving 1.03-1.11 higher MFU and training throughput.
 The performance gap becomes more pronounced as the model size decreases.
 This is because the smaller model size leads to a higher data parallelism (DP) size, which causes more inter-microbatch heterogeneity.
-In essence, MMScale’s data-aware reordering effectively mitigates data heterogeneity and enhances the training efficiency.
-(a) MFU. (b) Throughput.
+In essence, DistTrain’s disaggregated data reordering effectively mitigates data heterogeneity and enhances the training efficiency.
+We do not measure the running time of the reordering algorithm as it operates on dedicated CPU nodes asynchronously.
+It does not interfere with the GPU training process.
+The only overhead is the network delay which is evaluated in §8.3.
 | Model | # of | Global | Algorithm |
@@
 | MLLM-72B | 112 | 240 | 133ms |
-MMScale’s data-aware reordering algorithm executes before each training iteration.
-Table 3 presents a detailed analysis of the runtime overhead of the reordering algorithm.
-The DP size is the number of data parallelism groups in modality encoder.
-It completes the reordering in just a few milliseconds, a duration that is negligible compared to the several seconds to tens of seconds required for one iteration.
 8.3 Case Study
-In this subsection, we first evaluate MMScale under different frozen training settings. We then evaluate the effectiveness of system optimizations tailored for multimodal LLM training.
+In this subsection, we first evaluate DistTrain under different frozen training settings.
+We then evaluate the overhead of CPU preprocessing disaggregation and the effectiveness of StepCCL in mitigating the TP overhead.
 8.3.1 Frozen Training
@@
 All other experimental setup aligns with those detailed in §8.2.
-We evaluate the MFU and training throughput (TPT.) of MMScale compared to Megatron-LM.
+We evaluate the MFU and training throughput (TPT.) of DistTrain compared to Megatron-LM.
 The experimental results are presented in Figure 19 and Figure 20.
-MMScale consistently outperforms Megatron-LM across all frozen training configurations, achieving 1.4%-2.9% higher MFU and 1.2%-2.9% increased training throughput.
-This pronounced performance gap underscores the challenges posed by Megatron-LM’s rigid resource allocation in complex training environments.
-In contrast, MMScale adaptively adjusts resource allocation based on training settings and consistently achieves high resource utilization.
+DistTrain consistently outperforms Megatron-LM across all frozen training configurations, achieving 1.4-2.9 higher MFU and 1.2-2.9 higher training throughput.
+This pronounced performance gap underscores the challenges posed by Megatron-LM’s rigid model orchestration in complex training environments.
+In contrast, DistTrain adaptively adjusts model orchestration based on training settings and consistently achieves high resource utilization.
 (a) MFU. (b) Throughput.
-| Model | DP | Global | Intra-microbatch | Inter-microbatch |
-| Size | Batch Size | Reordering | Reordering | |
-| MLLM-72B | 30 | 1920 | 1.82ms | 5.87ms |
-| MLLM-72B | 60 | 1920 | 1.87ms | 3.02ms |
-| MLLM-72B | 120 | 1920 | 1.89ms | 1.71ms |
 (a) All modules freezing. (b) Exclusive encoder training. (c) Exclusive LLM training. (d) Exclusive generator training.
 (a) All modules freezing. (b) Exclusive encoder training. (c) Exclusive LLM training. (d) Exclusive generator training.
-7.3.2 System Optimizations
-Asynchronous Data Loader.
-We conduct an experiment to evaluate the effectiveness of asynchronous data loading optimization.
-Setting the DP size to one, we measure the average data loading time per iteration solely on the CPU.
-We then compare data loading times with and without this optimization and use varying numbers of images and different image resolutions for one training iteration.
-The results, depicted in Figure 21, indicate that asynchronous data loading significantly reduces loading time from seconds to milliseconds.
+Overhead of data preprocessing.
+We conduct an experiment to evaluate the overhead of data preprocessing, including decompression and reordering.
+Setting the DP size to one, we measure the average data preprocessing time per iteration on the GPU training side.
+We then compare data preprocessing time with and without CPU preprocessing disaggregation and use varying numbers of images and different image resolutions for one training iteration.
+The results, depicted in Figure 21, indicate that the disaggregation significantly reduces preprocessing time from seconds to milliseconds.
 The first parameter in the x-axis represents the number of images, while the second parameter denotes the image resolution.
-This reduction renders the loading time negligible relative to the total training time.
+In production training (§8.1), iteration times range from seconds to tens of seconds.
+Preprocessing overhead, initially counted in seconds, significantly interferes with training.
+With disaggregated data preprocessing, the overhead reduces to milliseconds, which is negligible relative to total iteration time.
 Mitigating TP overhead with StepCCL.
@@
 This assumption does not hold for multimodal LLMs due to the data heterogeneity in modality inputs of each sample.
-In contrast, MMScale leverages the specific training pattern of multimodal LLMs, creating a customized resource allocation problem that integrates tensor, pipeline, and data parallelism simultaneously.
-Besides, MMScale leverages data-aware reordering to harness the data heterogeneity.
+In contrast, DistTrain leverages the specific training pattern of multimodal LLMs, creating a customized model orchestration problem that integrates tensor, pipeline, and data parallelism simultaneously.
+Besides, DistTrain leverages disaggregated data reordering to harness the data heterogeneity.
 Sequence and expert parallelism.
@@
 These parallelism strategies are orthogonal to multimodal LLM training.
-In MMScale, both SP and EP are integrated into the training framework.
-MMScale treats SP and EP sizes as predefined parameters in the adaptive resource allocation optimization problem.
-Disaggregating multimodal modules from LLM.
-In MMScale, we address model heterogeneity by treating multimodal modules and LLM as separate stages in pipeline parallelism (PP).
-Another approach might integrate some LLM layers with multimodal modules into the same PP stage, while distributing the remaining LLM layers to maintain balance.
-However, this often becomes suboptimal since it requires multimodal modules to adopt the same TP strategy as the LLM layers within the same PP stage.
-Such alignment degrades performance when the optimal TP strategies for multimodal modules differ from those of LLM layers.
-Therefore, we choose to disaggregate multimodal modules from LLM as separate PP stages to achieve better performance.
+In DistTrain, both SP and EP are integrated into the training framework.
+DistTrain treats SP and EP sizes as predefined parameters in the disaggregated model orchestration optimization problem.
 (a) TP=4. (b) TP=8.
@@
 Efforts like [58, 59, 60] further optimize recovery process in cloud spot instance scenarios.
-These system optimizations of LLM training are orthogonal to MMScale.
+These system optimizations of LLM training are orthogonal to DistTrain.
 They overlook the model and data heterogeneity in multimodal LLMs.
-MMScale also integrates several of these optimizations in training the LLM backbone.
+DistTrain also integrates several of these optimizations in training the LLM backbone.
 Multimodal model training.
@@
 They fall short for scaling up to meet the demands of integrating multimodal models with LLMs which necessitate training across thousands of GPUs.
-This gap underpins the motivation behind MMScale, designed to meet the unique challenges of multimodal large language model training.
+This gap underpins the motivation behind DistTrain, designed to meet the unique challenges of multimodal large language model training.
 Multimodal LLM serving.
@@
 However, the serving of multimodal LLMs remains under-explored.
-MMScale’s core insights (e.g., disaggregated modules) are applicable to multimodal LLMs serving.
+DistTrain’s core insights (e.g., disaggregation) are applicable to multimodal LLMs serving.
 Our future work will delve into the specific challenges posed by multimodal LLMs serving systems.
 11 Conclusion
-We present MMScale, an efficient and adaptive framework to reform the training of multimodal LLMs.
+We present DistTrain, an efficient and adaptive framework to reform the training of multimodal LLMs.
 We identify the key challenges in training multimodal LLMs, i.e., model heterogeneity and data heterogeneity.
-MMScale introduces novel adaptive resource allocation to eliminate model heterogeneity and data-aware reordering to mitigate data heterogeneity.
-We evaluate MMScale on production cluster with thousands of GPUs and show that it achieves 54.7% MFU and outperforms Megatron-LM by up to 2.2 on throughput.
+DistTrain introduces disaggregated training to address these challenges, including disaggregated model orchestration to address model heterogeneity and disaggregated data reordering to address data heterogeneity.
+We evaluate DistTrain on production cluster with thousands of GPUs and show that it achieves 54.7% MFU and outperforms Megatron-LM by up to 2.2 on throughput.

### v3

@@
 DistTrain: Addressing Model and Data Heterogeneity with Disaggregated Training for Multimodal Large Language Models
-Abstract
-Multimodal large language models (LLMs) have demonstrated significant potential in a wide range of AI applications.
-Yet, training multimodal LLMs suffers from low efficiency and scalability, due to the inherent model heterogeneity and data heterogeneity across different modalities.
-We present DistTrain, an efficient and adaptive framework to reform the training of multimodal large language models on large-scale clusters.
-The core of DistTrain is the disaggregated training technique that exploits the characteristics of multimodal LLM training to achieve high efficiency and scalability.
-Specifically, it leverages disaggregated model orchestration and disaggregated data reordering to address model and data heterogeneity respectively.
-We also tailor system optimization for multimodal LLM training to overlap GPU communication and computation.
-We evaluate DistTrain across different sizes of multimodal LLMs on a large-scale production cluster with thousands of GPUs.
-The experimental results show that DistTrain achieves 54.7% Model FLOPs Utilization (MFU) when training a 72B multimodal LLM on 1172 GPUs and outperforms Megatron-LM by up to 2.2 on throughput.
-The ablation study shows the main techniques of DistTrain are both effective and lightweight.
-1 Introduction
-Recent advances in large language models (LLMs) are catalyzing a new wave of AI applications [1].
-However, LLMs are predominantly text-based, restricting their ability to understand and generate multimodal content.
-Emerging multimodal LLMs address this gap by integrating various modalities such as texts, images, and audios into LLMs which significantly enhances LLM’s applicability.
-Multimodal LLMs demonstrate great potential in tasks like image understanding [2, 3, 4, 5], audio comprehension [6, 7], and embodied AI [8, 9].
-Many organizations are actively developing their multimodal LLMs, such as OpenAI’s GPT-4o [10], Google’s Gemini [11] and PaLM-E [9], Meta’s Chameleon [12], etc.
-Training multimodal LLMs demands vast computational resources.
-According to the scaling law [13], model size and training data volume are crucial for determining model capabilities.
-Substantial efforts are invested in training models with billions of parameters on trillion-scale tokens.
-For instance, Meta’s Chameleon [12] is a 34B multimodal LLM trained on more than 4.8 trillion tokens.
-Leading organizations often deploy large-scale clusters, equipped with thousands of GPUs, for such training tasks.
-One technical report [14] reveals that training a GPT-3 175B with 300 billion tokens costs $4.6 million and lasts one year using 355 V100 GPUs.
-Consequently, it is crucial to develop an efficient and scalable training framework to minimize costs and accelerate training.
-Figure 1 depicts the mainstream model architecture of multimodal LLMs, comprising three primary modules: the modality encoder, LLM backbone, and the modality generator [15, 16].
-These modules are linked by the projector, which may incorporate MLP or cross-attention layers.
-The modality encoder transforms input from various modalities into a unified embedding space.
-The embedding vectors are then analyzed by the LLM backbone, a transformer model, to discern data patterns and inter-modal relationships.
-Subsequently, the modality generator translates this processed information back into coherent outputs tailored to each modality.
-Existing LLM training framework, e.g., Megatron-LM [17], can be extended to train multimodal LLMs by treating the multimodal modules as additional layers within the LLM.
-However, training multimodal LLMs poses two substantial challenges: model heterogeneity and data heterogeneity.
-The fundamental issue of model heterogeneity stems from the need to process diverse modalities with different modules that vary dramatically in size and operator complexity.
-The model heterogeneity across different modules (i.e., modality encoder, LLM backbone, and modality generator) introduces severe pipeline bubbles, resulting in poor GPU utilization.
-Meanwhile, data heterogeneity bursts onto the scene due to the intricate and unstructured nature of multimodal input data.
-The data heterogeneity across different modality data leads to inter-microbatch and intra-microbatch training stragglers, which prolong the training duration and exacerbate the pipeline bubbles.
-These challenges collectively limit the efficiency and scalability of multimodal LLM training, resulting in MFU as low as ~20% in production-level training (§8.1).
-To this end, we present DistTrain, an efficient and adaptive framework to reform the training of multimodal LLMs.
-DistTrain achieves state-of-the-art MFU which is close to unimodal LLM training and effectively scales to large clusters with thousands of GPUs.
-The core principle of DistTrain is disaggregated training, including GPU training disaggregation and CPU preprocessing disaggregation, which facilitate to address model and data heterogeneity respectively.
-For model heterogeneity, we meticulously analyze the pipeline bubbles stemming from model heterogeneity and identify their root causes (§2.2).
-The GPU training disaggregation of modality encoder, LLM backbone, and modality generator enables adaptive orchestration across the three modules in multimodal LLM.
-Building on this, we propose disaggregated model orchestration that navigates the complicated design space to choose the optimal resource and parallelism configurations.
-This innovative approach minimizes the pipeline bubbles caused by model heterogeneity and achieves optimal training efficiency.
-For data heterogeneity, we categorize the training stragglers into inter-microbatch and intra-microbatch stragglers (§2.3).
-The CPU preprocessing disaggregation allows efficient data preprocessing (e.g., data decompression and reordering) with negligible runtime overhead.
-From this foundation, we incorporate disaggregated data reordering into the preprocessing to strategically reorder training data without additional overhead.
-The inter-microbatch reordering algorithm reorders the data samples to evenly distribute the load across different data parallelism groups.
-The intra-microbatch reordering algorithm reorders the microbatches tailored for (interleaved) 1F1B pipeline scheme to minimize the pipeline bubbles.
-This two-level approach effectively mitigates data heterogeneity.
-In addition, we customize system optimization for multimodal LLM training.
-We implement an in-house collective communication library, StepCCL, to hide the communication overhead within the computation.
+Abstract.
+Multimodal large language models (LLMs) empower LLMs to ingest inputs and generate outputs in multiple forms, such as text, image, and audio.
+However, the integration of multiple modalities introduces heterogeneity in both the model and training data, creating unique systems challenges.
+We propose DistTrain, a disaggregated training system for multimodal LLMs.
+DistTrain incorporates two novel disaggregation techniques to address model and data heterogeneity, respectively.
+The first is disaggregated model orchestration, which separates the training for modality encoder, LLM backbone, and modality generator.
+This allows the three components to adaptively and independently orchestrate their resources and parallelism configurations.
+The second is disaggregated data preprocessing, which decouples data preprocessing from training.
+This eliminates resource contention between preprocessing and training, and enables efficient data reordering to mitigate stragglers within and between microbatches caused by data heterogeneity.
+We evaluate DistTrain across different sizes of multimodal LLMs on a large-scale production cluster.
+The experimental results show that DistTrain achieves 54.7% Model FLOPs Utilization (MFU) when training a 72B multimodal LLM on 1172 GPUs and outperforms Megatron-LM by up to 2.2 on training throughput.
+ACM Reference Format:
+Zili Zhang, Yinmin Zhong, Yimin Jiang, Hanpeng Hu, Jianjian Sun, Zheng Ge, Yibo Zhu, Daxin Jiang, Xin Jin.
+2025.
+DistTrain: Addressing Model and Data Heterogeneity with Disaggregated Training for Multimodal Large Language Models.
+In SIGCOMM ’25, August 8-August 11, 2025, Coimbra, Portugal, 15 pages.
+DistTrain: Addressing Model and Data Heterogeneity with Disaggregated Training …
+1. Introduction
+Traditional multimodal models, which focus solely on either cross-modal feature representation (e.g., CLIP (Radford et al., 2021)) or multimodal content generation (e.g., Stable-Diffusion (Rombach et al., 2022) and VAE (Kingma et al., 2013)), cannot simultaneously process multimodal inputs and generate multimodal outputs.
+This restricts their capabilities and hinders seamless human interaction.
+Multimodal large language models (LLMs) emerge to integrate the strengths of LLMs (Vaswani et al., 2017; Brown et al., 2020; Touvron et al., 2023) with traditional multimodal models to simultaneously ingest multimodal inputs and generate multimodal outputs in auto-regressive manner.
+Multimodal LLMs have already shown great potential in tasks such as vision understanding and generation (Liu et al., 2024b; Zhang et al., 2023b; Wang et al., 2024; Dong et al., 2023; Kondratyuk et al., 2023), audio comprehension (Rubenstein et al., 2023; Borsos et al., 2023), and embodied AI (Zhang et al., 2023a; Driess et al., 2023).
+Many organizations are actively developing multimodal LLMs, such as OpenAI’s GPT-4o (gpt, 2024), Google’s Gemini (gem, 2024), and Meta’s Llama3 (lla, 2024), etc.
+Figure 1 depicts the model architecture of multimodal LLMs, comprising three modules: modality encoder, LLM backbone, and modality generator (Dong et al., 2023; Yin et al., 2023; Zhang et al., 2024).
+We emphasize this architecture underpins most state-of-the-art multimodal LLMs (Table 1).
+These modules are linked by projectors, which may incorporate MLP or cross-attention.
+The modality encoder transforms inputs from various modalities into different modality tokens.
+These tokens from different modality are interleaved as a sequence for LLM training.
+The modality generator translates the processed information back into coherent outputs tailored to each modality.
+Training multimodal LLMs is more challenging than training traditional multimodal models.
+Traditional multimodal model training frameworks like DistMM (Huang et al., 2024) lack support for simultaneous training of encoders, LLMs, and generators, particularly techniques tailored for LLMs.
+The current practice for training multimodal LLMs (Alayrac et al., 2022; Wu et al., 2024) extends LLM training frameworks (e.g., Megatron-LM (Shoeybi et al., 2019)) with additional modality encoder and generator modules, which uses a monolithic approach.
+In this context, the monolithic training refers to two aspects.
+First, it uses the same data and tensor parallelism strategies of distributed training across the modality encoder, LLM backbone, and modality generator modules.
+As for pipeline parallelism, the encoder and generator modules are incorporated as additional pipeline stages in the training pipeline.
+Second, it couples data preprocessing into training process, and co-locates data preprocessing and training on the same machines.
+The monolithic approach falls short of handling unique systems challenges in multimodal LLM training.
+Compared to traditional LLM training that only handles text, the multiple modalities in multimodal LLM training introduces model heterogeneity and data heterogeneity.
+Model heterogeneity stems from processing multiple modalities with different modules (i.e., encoder, LLM backbone, and generator) that vary dramatically in size and operator complexity.
+Using the same parallelism strategies across different modules introduces severe pipeline bubbles, resulting in poor GPU utilization (§2.2).
+Data heterogeneity emerges from processing multimodal data that differ significantly in volume and processing cost, e.g., a high-resolution image is much larger and requires more resources for preprocessing than a line of text.
+In the monolithic approach, where data preprocessing and training are integrated and co-located, the substantial processing cost significantly hinders the training process.
+Additionally, the heterogeneity across different modality data (i.e., each training sample has a varying number of modality tokens) leads to inter-microbatch and intra-microbatch training stragglers, which prolong the training duration and exacerbate the pipeline bubbles (§2.3).
+In our initial efforts using the monolithic approach, we observed extremely low MFU in production training, as low as 20% (§7.1).
+To this end, we propose DistTrain, a disaggregated training system for multimodal LLMs.
+DistTrain applies the systems principle of disaggregation in the context of multimodal LLM training to address the unique systems challenges.
+We design two novel disaggregation techniques, which are disaggregated model orchestration and disaggregated data preprocessing, to address model and data heterogeneity, respectively.
+For model heterogeneity, we analyze the pipeline bubbles caused by multiple modalities across different modules in multimodal LLMs, and identify their root causes (§2.2).
+To address this, we design disaggregated model orchestration to separate the training of modality encoder, LLM backbone, and modality generator.
+This separation enables adaptive orchestration of resources and parallelism configurations across the three modules.
+Building on this, we formulate the model orchestration as an optimization problem aimed at minimizing the training time per iteration.
+We then design an adaptive model orchestration algorithm that navigates the complicated search space to find the optimal resource and parallelism configurations.
+Disaggregated model orchestration is able to minimize pipeline bubbles caused by model heterogeneity and achieves optimal training efficiency.
+For data heterogeneity, we characterize the training data of different modalities, and unveil their preprocessing overheads and the stragglers they brought into the training process (§2.3).
+We categorize the stragglers into intra-microbatch and inter-microbatch stragglers.
+We design disaggregated data preprocessing to eliminate the interference between preprocessing and training.
+Importantly, such disaggregation enables flexible data preprocessing, and provides new opportunities for data reordering to mitigate stragglers without additional overhead.
+We use intra-microbatch reordering to evenly distribute the load across data parallelism groups, and inter-microbatch reordering to hide the training stragglers into the training pipeline.
+This dual-level reordering effectively mitigates the stragglers caused by data heterogeneity.
+Importantly, since the reordering only permutes the sequence of commutative gradient accumulation operations, it preserves the training’s convergence semantics.
 In summary, we make the following contributions.
-We present DistTrain, an efficient and adaptive framework to reform the training of multimodal LLMs by addressing the multimodal heterogeneity.
-It delivers state-of-the-art MFU on large-scale clusters with thousands of GPUs.
-We identify and discuss the primary challenges associated with multimodal LLM training, which are summarized as model heterogeneity and data heterogeneity.
-We propose disaggregated training for multimodal LLM.
-It leverages disaggregated model orchestration and disaggregated data reordering to effectively address model and data heterogeneity, respectively.
+We identify the unique system challenges introduced by multiple modalities in multimodal LLM training, and summarize them as model and data heterogeneity.
+We propose DistTrain, a disaggregated training system for multimodal LLMs, and design two novel techniques—disaggregated model orchestration and disaggregated data preprocessing—to address model and data heterogeneity.
 We implement DistTrain and conduct experiments on our production cluster with thousands of GPUs.
-The experimental results show that DistTrain achieves 54.7% MFU when training a 72B multimodal LLM on 1172 GPUs and outperforms Megatron-LM by up to 2.2 on throughput.
-2 Motivation
-2.1 Multimodal Large Language Model Training
-Large language model.
-Large language models (LLMs) [18, 19, 20] have revolutionized natural language processing (NLP) by achieving state-of-the-art performance on a wide range of tasks, such as text generation, translation, and summarization.
-Many organizations have raced to develop their own LLMs, such as OpenAI’s GPT-4 [21], Google’s Gemini [11], and Meta’s Llama [22].
-The core architecture of LLMs consists of a stack of homogeneous transformer layers [18] that use self-attention mechanisms to capture contextual information in text.
-LLMs are pre-trained with unsupervised learning on large-scale text corpora and then fine-tuned on task-specific text datasets.
-The text data is typically tokenized into fixed-length sequences, which are then fed into the model to learn the underlying patterns.
-The training process involves weeks or even months of computation on dedicated AI clusters with thousands of GPUs.
-According to one technical report [14], training 175B GPT-3 requires 4.6 million dollars.
-Optimizing the training process is essential to reduce the stupendous cost and accelerate the model deployment.
-Multimodal LLM.
-The unimodal LLMs are limited to processing text data, which restricts their applicability to multimodal tasks (e.g., image understanding and generation).
-As a result, multimodal LLMs have emerged to address this limitation by integrating multiple modalities (e.g., images, audios, and videos) into the advanced LLMs [15], which support multimodal inputs and outputs during LLM generation.
-For example, GPT-4o [10] garners widespread attention by facilitating more natural interactions with humans through both visual and auditory modalities.
-Moreover, the predominantly text-based data in human societies is finite [23].
-Harnessing multimodal data is inevitable to continually expand and enhance LLM capabilities.
-Figure 1 illustrates the model architecture of a multimodal LLM, which consists of three modules: a modality encoder, an LLM backbone, and a modality generator [24, 15, 16].
-The modality encoder transforms input data from different modalities (e.g., ViT [25] for images and Beats [26] for audios) into an intermediate representation (i.e., an embedding tensor), which is then projected into a unified embedding space across modalities with input projection layers (e.g., MLP and cross-attention).
-The LLM backbone, typically a transformer model (e.g., GPT [19, 20] and Llama [22]), processes the multimodal embeddings to discern the intricate data patterns and inter-modal relationships.
-The output data of the LLM backbone is subsequently refined by output projection layers, which tailor the information for each modality.
-Finally, the modality generator (e.g., Diffusion [27] for images and AudioLDM [28] for audios) transforms the LLM-processed information back into the respective modal outputs for generation.
+The experimental results show that DistTrain achieves 54.7% MFU when training a 72B multimodal LLM on 1172 GPUs and outperforms Megatron-LM by up to 2.2 on training throughput.
+2. Motivation
+2.1. Multimodal LLM Training
+Traditional multimodal models.
+Traditional multimodal models encompass diverse architectures, such as contrastive learning (Radford et al., 2021) and stable diffusion (Rombach et al., 2022) (SD).
+Contrastive learning uses encoder models to capture cross-modal similarities, primarily for feature representation.
+Stable diffusion generates images from text prompts by iteratively refining random noise through a denoising process, enabling text-to-image generation.
+However, these models are task-specific and lack the flexibility to handle diverse multimodal scenarios, which cannot process multimodal inputs and generate multimodal outputs simultaneously, restricting their model capabilities in complex real-world settings and hindering seamless interaction with humans.
+Multimodal | Encoder(s) | LLM | Generator(s) |
+LLM | Backbone | ||
+Flamingo (Alayrac et al., 2022) | NFNet (Brock et al., 2021) | GPT-3 (Brown et al., 2020) | LM-Head |
+LLaVA (Liu et al., 2024a) | CLIP (Radford et al., 2021) | Vicuna (chi, 2023) | LM-Head |
+PaLM-E (Driess et al., 2023) | ViT (Dosovitskiy, 2020) | PaLM (Chowdhery et al., 2023) | LM-Head |
+EMU (Sun et al., 2023b) | EVA-CLIP (Sun et al., 2023a) | Llama (Touvron et al., 2023) | LM-Head, SD (Rombach et al., 2022) |
+Bagel (Deng et al., 2025) | ViT (Sun et al., 2023a) | Qwen2.5 (yan, 2024) | LM-Head, VAE (Kingma et al., 2013) |
+VideoPoet (Kondratyuk et al., 2023) | MAGViT (Yu et al., 2023), | GPT (Radford et al., 2018) | MAGViT (Yu et al., 2023), |
+SoundStream (Zeghidour et al., 2021) | SoundStream (Zeghidour et al., 2021) |
+Multimodal LLMs.
+Large language models (LLMs)(Achiam et al., 2023; gem, 2024; lla, 2024), built on homogeneous transformer layers and trained on extensive text corpora, excel at understanding and generating high-quality text but are limited to text-only processing.
+Multimodal LLMs address these limitations by integrating the strengths of LLMs with traditional multimodal models, enabling the simultaneous processing of diverse inputs (e.g., images, audio, video) and generating multimodal outputs in auto-regressive manner.
+For example, GPT-4o (gpt, 2024) gains wide attention by facilitating more natural interaction with humans through both visual and auditory modalities.
+Figure 1 illustrates the model architecture of a multimodal LLM, which consists of three modules: a modality encoder, an LLM backbone, and a modality generator (Alayrac et al., 2022; Yin et al., 2023; Zhang et al., 2024).
+Different from traditional multimodal models, this generative model facilitates a more robust and interpretable process for understanding and generating multimodal contents through auto-regressive decoding.
+Specifically, the modality encoder transforms input data from different modalities (e.g., ViT (Dosovitskiy, 2020) for images and Beats (Chen et al., 2022) for audios) into an intermediate representation, which is then converted into different modality tokens with input projector (e.g., MLP and cross-attention).
+These tokens from different modalities form a sequence for LLM training.
+The LLM backbone, typically a transformer model (e.g., GPT (Radford et al., 2019; Brown et al., 2020) and Llama (lla, 2024)), processes the input sequence to extract intricate data patterns and inter-modal relationships.
+The LLM backbone’s output is refined by an output projector and then passed to a modality generator (e.g., Diffusion (Rombach et al., 2022) for images, AudioLDM (Liu et al., 2023a) for audio), which converts the processed information into the corresponding modal outputs.
+Table 1 summarizes the three modules of some state-of-the-art multimodal LLMs.
+We omit text tokenizer (e.g., BPE) and encoder (e.g., text embedding) from the table’s encoder column.
+LM-Head is responsible for decoding the text tokens
 Multimodal LLM training necessitates training all three modules simultaneously.
-Additionally, during different training phases, specific modules are frozen to stabilize training loss and enhance model effectiveness [15].
-Regarding the training data, the input sequence comprises text, image, and audio tokens.
-These data from different modalities are tokenized into subsequences which are then interleaved to form fixed-length training sequences [12].
-In cases where the input sequence falls short, it is padded with zero tokens for batching.
-Distinct modality sub-sequences are demarcated by special tokens and processed through modality-specific encoders.
-Training framework.
-To train the multimodal large language model with large-scale clusters, the de facto solution is to leverage Megatron-LM, a highly efficient and robust training framework for large-scale transformer models.
-Megatron-LM employs a unified parallelism strategy for the entire model.
-It combines tensor parallelism (TP) and pipeline parallelism (PP) to distribute the model parameters across multiple GPUs.
-TP divides the model parameter within each layer, while PP partitions parameters between layers.
-It also leverages data parallelism (DP) to distribute the training data.
-For multimodal LLMs, Megatron-LM is easily extended to integrate additional multimodal modules.
-Figure 2 illustrates the training topology of multimodal LLMs with Megatron-LM.
-Specifically, Megatron-LM treats the multimodal modules as additional layers within the LLM and incorporates additional PP stages to accommodate the modality encoder and generator.
-The same TP strategy used in the LLM backbone is applied to these two multimodal modules.
-If the modality encoder and generator are not large enough, they are replicated across the GPUs in the TP group to maximize resource utilization.
-As for DP, Megatron-LM applies the same DP strategy to the multimodal modules as the LLM backbone.
-The projector is co-located with the modality encoder and generator and is replicated across the GPUs in the TP group.
-However, this training framework introduces significant computation imbalance stemming from model heterogeneity and data heterogeneity due to its rigid model orchestration method.
-(i.e., the multimodal modules share the same DP and TP strategies with LLM backbone).
-It only achieves ~20% MFU (§8.1) which is significantly lower than the MFU of ~50% observed in training unimodal (text-only) LLMs [29].
-2.2 Model Heterogeneity
-(a) Distribution of text subsequence size. (b) Distribution of image subsequence size. (c) Distribution of image subsequence count.
-The first challenge is the computation imbalance arising from model heterogeneity.
-Each module in multimodal LLMs bears different computational demands due to varying operators and inputs.
-For instance, ViT, as modality encoder, is constructed with narrow transformer layers (i.e., small hidden size), whereas LLM backbone is built on broader transformer layers (i.e., large hidden size).
-Meanwhile, Diffusion, as modality generator, utilizes a combination of convolution and attention layers (i.e., U-Net).
-This architectural diversity results in distinct computation time for each module.
+Additionally, during different training phases, specific modules are frozen to stabilize training loss (Chen et al., 2025; Deng et al., 2025).
+Regarding the data, input sequences comprise text, image, and audio tokens.
+The data from different modalities are encoded into subsequences which are then interleaved to form fixed-length training sequences (Team, 2024).
+Training frameworks.
+Multimodal model training frameworks such as DistMM (Huang et al., 2024) do not support training encoder, LLM backbone and generator simultaneously.
+Especially, they lack support training techniques (Chen et al., 2024; Narayanan et al., 2021; Rajbhandari et al., 2020) for LLMs.
+In contrast, LLM training frameworks like Megatron-LM (Shoeybi et al., 2019) can be extended with modality encoder and generator modules, making them widely used in production multimodal LLM training, e.g., Flamingo (Alayrac et al., 2022), DeepSeek-VL2 (Wu et al., 2024) and VideoPoet (Kondratyuk et al., 2023).
+Figure 2 shows training multimodal LLMs with Megatron-LM.
+Megatron-LM adds multimodal modules as additional layers, and incorporates additional pipeline parallelism (PP) stages to accommodate encoder and generator.
+The same tensor parallelism (TP) strategy used in the LLM backbone is applied to encoder and generator.
+If the encoder and generator are not large enough, they are replicated across the GPUs in the TP group to maximize resource utilization.
+As for data parallelism (DP), Megatron-LM applies the same DP strategy to the multimodal modules as the LLM backbone.
+The projector is co-located with the encoder and generator, and is replicated across the GPUs in the TP group.
+Moreover, data preprocessing is incorporated into the training process and co-located with training on the same node.
+This approach is monolithic in terms that the encoder and generator use the same parallelism strategy as the LLM backbone, and that preprocessing and training are integrated and co-located.
+Such a monolithic approach introduces significant computation imbalance stemming from model and data heterogeneity.
+2.2. Model Heterogeneity
+Characterization.
+Each module in multimodal LLMs has different computational demands due to varying operators and inputs.
+For instance, ViT, as modality encoder, is constructed with narrow transformer layers (i.e., small hidden size), whereas the LLM backbone is built with wide transformer layers (i.e., large hidden size).
+Meanwhile, Diffusion, as modality generator, utilizes a combination of convolution and attention layers.
+The diversity in model architecture results in distinct computation time for each module.
 Figure 3 shows varying forward time under different input configurations with Megatron-LM.
@@
 The first configuration parameter is the number of images in the 8K input sequence, and the second is the image resolution.
-The time differs markedly across different configurations.
-The computational imbalance between modules leads to two types of pipeline bubbles in pipeline parallelism.
-The first arises in the modality encoder and generator stages, as shown in Figure 4(a), resulting from their inadequate utilization of assigned GPU resources.
-The second type emerges in the stages of the LLM backbone, as shown in Figure 4(b).
-This is because the intensive computational demands of the encoder and generator extend their stage durations.
+The results reveal a key performance disparity: while the LLM backbone maintains a consistent forward time, the modality encoder and generator exhibit significant variations.
+(a) Distribution of text subsequence size. (b) Distribution of image subsequence size. (c) Distribution of image subsequence count.
+Pipeline imbalance.
+The computation imbalance between modules leads to two types of bubbles in pipeline parallelism.
+The first arises in the modality encoder and generator stages, as shown in Figure 4(a), resulting from their inadequate utilization of allocated GPU resources.
+The second type arises in the stages of the LLM backbone, as shown in Figure 4(b).
+This is because the intensive computation demands of the encoder and generator increase their stage durations.
 Due to the pipeline dependency, the LLM stages are forced to wait for the multimodal stage to complete, thereby creating pipeline bubbles.
 The latter problem is particularly pronounced during large-scale multimodal LLM training, where the bulk of GPU resources are allocated to the LLM backbone.
-These pipeline bubbles, stemming from model heterogeneity, substantially diminish the MFU during the training.
-2.3 Data Heterogeneity
-The second challenge is the computational imbalance stemming from data heterogeneity.
-Each input sequence (i.e., training sample) for multimodal LLM training consists of interleaved modality subsequences that exhibit highly skewed distributions.
-Focusing on images and texts, we perform data characterization on the LAION-400M dataset [30], an open-source collection of images paired with text captions.
-Each image (i.e., one image subsequence) is segmented into 1616 patches, and each patch is tokenized into one image token.
+These pipeline bubbles, stemming from model heterogeneity, reduce the MFU significantly during training.
+2.3. Data Heterogeneity
+Characterization.
+The training data samples of Multimodal LLMs often combine lightweight text with heavyweight multimodal data.
+The latter significantly increases data preprocessing time.
+For example, a typical training sample could include a 256-word text sequence and ten 10241024 RGB images.
+The text is just kilobytes, whereas the images are total of 120 megabytes.
+Preprocessing (e.g., decompression, resizing, and reordering) such samples can take several seconds, and prolong the co-located training process.
+Additionally, each input sequence consists of interleaved modality subsequences that exhibit highly skewed distributions.
+Focusing on images and texts, we perform data characterization on the LAION-400M dataset (Schuhmann et al., 2021).
+Each image (i.e., one image subsequence) is segmented into 1616 patches, and each patch is converted into one image token.
 The texts are tokenized through Llama tokenizer.
-The image tokens are interleaved with text tokens to create an 8K-token input sequence for training.
-As shown in Figure 5(a) and Figure 5(b), the sizes of text and image subsequences display distinctly skewed distributions.
+The image tokens are interleaved with text tokens to create an 8K-token training sequence.
+As shown in Figure 5(a) and Figure 5(b), the sizes of text and image subsequences display highly skewed distributions.
 We further analyze the count of modality subsequence per training sample using image as an example.
-The count of image subsequences per training sample, shown in Figure 5(c), also demonstrates a skewed distribution.
-Different sample size (i.e., modality tokens per sample) leads to varying computation time in the modality encoder and generator stages.
-Such data heterogeneity results in both intra-microbatch and inter-microbatch stragglers within the pipeline parallelism (PP) stages of the modality encoder and generator.
-These stragglers exacerbate the computational imbalances and further reduce the GPU utilization.
-It is noted that all microbatches within LLM backbone have the same computation time since the sequence length is fixed.
-We do not consider data heterogeneity between global batches, as each global batch contains numerous randomly shuffled training samples (e.g., thousands with a large DP size), which effectively smooths out the data heterogeneity.
-Intra-microbatch straggler.
-Intra-microbatch straggler occurs when particularly large training samples decelerate training, as DP groups handle variably-sized training samples.
+The count of image subsequences per training sample, shown in Figure 5(c), also shows a skewed distribution.
+Different sample sizes (i.e., modality tokens per sample) lead to varying computation time in the modality encoder and generator.
+Stragglers.
+Data heterogeneity results in intra-microbatch and inter-microbatch stragglers within the PP stages of the modality encoder and generator.
+These stragglers exacerbate the computation imbalances and further reduce GPU utilization.
+It is noted that all microbatches within the LLM have the same computation time since the sequence length is fixed.
+We do not consider data heterogeneity between global batches, as each global batch contains numerous randomly shuffled training samples (e.g., thousands with a large DP size), which effectively smooths out data heterogeneity.
+Intra-microbatch straggler arises since different DP groups handle variably-sized training samples.
 Illustrated in Figure 6, the first DP group () processes two large training samples within two microbatches.
-In contrast, the second DP group () processes two smaller samples in the same microbatches, completing them more swiftly.
+In contrast, the second DP group () processes two smaller samples in the same microbatches, completing them more quickly.
 Consequently, lags behind and becomes the straggler, which delays the overall training process.
-Inter-microbatch straggler.
 Inter-microbatch straggler emerges from pipeline imbalances between microbatches.
 As depicted in Figure 7, the first pipeline stage is the modality encoder followed by one LLM backbone stage.
-Figure 7(a) illustrates the pipeline without data heterogeneity, where the modality encoder processes each microbatch with consistent time.
-In contrast, Figure 7(b) depicts the pipeline with data heterogeneity, where the forward time of the modality encoder varies markedly across microbatches.
-The straggler (i.e., the microbatch ) significantly delays the training process of the subsequent PP stages, leading to a large pipeline bubble.
-3 DistTrain Overview
-We present DistTrain, an efficient and adaptive framework to reform the training of multimodal LLMs by addressing the multimodal heterogeneity.
-DistTrain proposes disaggregated training for multimodal LLM to achieve high efficiency and scalability.
-DistTrain eliminates the model heterogeneity by disaggregated model orchestration (§5) and harnesses the data heterogeneity by disaggregated data reordering (§6).
-In addition, DistTrain adopts some system optimizations tailored for multimodal LLM training (§7).
-Here we provide a brief overview of DistTrain as Figure 8 shows.
+Figure 7(a) illustrates the pipeline without data heterogeneity, where the modality encoder processes each microbatch with the same amount of time.
+In contrast, Figure 7(b) depicts the pipeline with data heterogeneity, where the forward time of the modality encoder varies largely across microbatches.
+The straggler (i.e., the microbatch a) significantly delays the training process of the subsequent PP stages, leading to a large pipeline bubble.
+3. DistTrain Overview
+We propose DistTrain, a disaggregated training system for multimodal LLMs.
+DistTrain addresses model heterogeneity by disaggregated model orchestration (§4) and handles data heterogeneity by disaggregated data preprocessing (§5).
+Figure 8 shows an overview of DistTrain.
 DistTrain manager.
 Before training, DistTrain employs a training manager to determine the resource allocation and parallelism strategy for each module in multimodal LLMs.
-This scheduler first gathers the model architecture and training configuration (e.g., global batch size) from the user and randomly samples a subset of training data to analyze the data distribution.
-Utilizing this information, it runs a series of benchmarking training trials and constructs a performance profiler with linear interpolation to estimate each module’s computation and communication time.
+The training manager first gathers the model architecture and training configuration (e.g., global batch size) from the user and samples a subset of training data to analyze the data distribution.
+Utilizing the information, it runs a series of benchmarking training trials and constructs a performance profiler with linear interpolation to estimate each module’s computation and communication time.
 Based on the profiling results, the training manager decides the optimal resource allocation and parallelism strategy with disaggregated model orchestration for one specific training task, as detailed in §4.
 DistTrain initializer.
-DistTrain then initializes the GPU training disaggregation for modality encoder, LLM backbone, and modality generator, respectively.
-DistTrain allocates different numbers of GPUs to each parallelism unit.
-Each unit then establishes the specific communication group.
-The unit loads the model checkpoint from distributed file system and shards the model parameters and optimization states.
+DistTrain then initializes the disaggregated model orchestration for modality encoder, LLM backbone, and modality generator, respectively.
+DistTrain allocates different numbers of GPUs and determines the parallelism strategy for each module.
+Each module then establishes the specific communication group.
+It then loads the model checkpoint from a distributed file system and shards the model parameters.
 Finally, DistTrain conducts several communication trials to warm up the system and test connectivity.
 DistTrain runtime.
-At runtime, the dedicated CPU nodes (i.e., CPU preprocessing disaggregation) retrieve training samples from the distributed file system for preprocessing.
-It performs disaggregated data reordering to reorder the training samples within one global batch without breaking the synchronous training semantics [31].
-This reordering effectively eliminates both inter-microbatch and intra-microbatch data heterogeneity, as detailed in §6.
-In each iteration, the main training process receives the preprocessed data asynchronously from the CPU nodes.
+At runtime, the dedicated CPU nodes (i.e., disaggregated data preprocessing nodes) retrieve training data samples from the distributed file system for preprocessing.
+It performs data reordering to reorder the training samples within one global batch while preserving the synchronous training semantics (alp, 2022).
+The reordering mitigates both intra-microbatch and inter-microbatch stragglers caused by data heterogeneity, as detailed in §5.
+In each iteration, the GPU training nodes receive the preprocessed data asynchronously from the CPU nodes.
 The data then undergoes sequentially through the modality encoder, LLM backbone, and modality generator in the training pipeline.
-Finally, it synchronizes the gradients and model parameters through the all-gather operation, employing the ZERO-1 optimization [32] and mixed precision training [33].
-Additionally, DistTrain adopts a dedicated process to periodically and asynchronously save the model checkpoint to the distributed file system for fault tolerance.
-4 Disaggregated Training
-To address the model and data heterogeneity in multimodal LLM training, we first introduce DistTrain’s core principle: disaggregated training.
-It includes GPU training disaggregation and CPU preprocessing disaggregation.
-GPU training disaggregation provides opportunities for adaptive model orchestration across the three modules to address the model heterogeneity.
-CPU preprocessing disaggregation facilitates data preprocessing to address data heterogeneity with negligible runtime overhead.
-4.1 GPU Training Disaggregation
-Figure 9 demonstrates the training topology of disaggregated GPU training in DistTrain.
-Different from the rigid model orchestration in Megatron-LM (i.e., Figure 2), DistTrain is able to adaptively adjust the resource allocation and parallelism strategy.
-For instance, DistTrain allocates 4 GPUs (DP=2 and TP=2) to the modality encoder, 12 GPUs (DP=3 and TP=4) to the LLM backbone per PP stage, and 4 GPUs (DP=1 and TP=4) to the modality generator.
+Finally, the GPU training nodes synchronize the gradients and model parameters through collective communication.
+DistTrain employs ZERO-1 optimization (Rajbhandari et al., 2020) and mixed precision training (Micikevicius et al., 2017) to reduce the memory footprint and accelerate the training of LLM backbone.
+Additionally, DistTrain adopts a dedicated process to periodically and asynchronously save model checkpoints to the distributed file system for fault tolerance.
+4. Addressing Model Heterogeneity
+To address model heterogeneity, we first introduce disaggregated model orchestration to adaptively orchestrate the resource and parallelism configurations of the three modules in multimodal LLM training.
+We then formulate the problem to minimize the training time per iteration.
+Last, we present the adaptive model orchestration algorithm to find optimal resource and parallelism configurations.
+4.1. Disaggregated Model Orchestration
+Figure 9 illustrates the disaggregated model orchestration.
+Different from the monolithic model orchestration in Megatron-LM (i.e., Figure 2), DistTrain is able to adaptively adjust the resource allocation and parallelism strategy.
+For instance, DistTrain can allocate 4 GPUs (DP=2 and TP=2) to the modality encoder, 12 GPUs (DP=3 and TP=4) to the LLM backbone per PP stage, and 4 GPUs (DP=1 and TP=4) to the modality generator.
 Additionally, the projector layers are co-located with either the modality encoder or generator, with their number of replicas adapting as needed.
-We implement GPU training disaggregation through a dedicated module, i.e., parallelism unit.
+We implement disaggregated model orchestration through a dedicated module, i.e., parallelism unit.
 Parallelism unit.
@@
 DistTrain treats the modality encoder, LLM backbone, and modality generator as three individual parallelism units.
-The detailed implementation of parallelism unit is discussed in §7.
-4.2 CPU Preprocessing Disaggregation
-When training multimodal LLMs, training samples often combine lightweight text with heavyweight multimodal data.
-The latter significantly increases data preprocessing time.
-For example, a typical training sample could include a 256-word text sequence and ten 10241024 RGB images.
-The text is just kilobytes, whereas the images are total of 120 megabytes.
-Preprocessing (e.g., decompression, resizing, and reordering) such samples can take several seconds and interfere with the co-located training process.
-DistTrain disaggregates the CPU data preprocessing from the GPU training process with a producer-consumer model.
-The producer, operating on dedicated CPU nodes, fetches data from the distributed file system and preprocesses training data asynchronously with the GPU training process.
-The consumer, i.e., the main training process, receives this preprocessed data for training.
-The producer and consumer communicate through RPC calls, and use RDMA network for lower latency if available.
-This disaggregation guarantees that the CPU preprocessing does not interfere with the GPU training process and achieves negligible data preprocessing overhead.
-5 Addressing Model Heterogeneity
-Disaggregated training enables disaggregated model orchestration among the different modules based on the training workload.
-We first formulate the problem of disaggregated model orchestration to minimize the training time per iteration.
-Then, we present the detailed algorithm to optimally address the problem caused by model heterogeneity.
-5.1 Problem Formulation
-With disaggregated training, we are able to adaptively orchestrate the three modules.
+For advanced workloads, DistTrain extends the capabilities of these units.
+To handle long sequences, it integrates sequence parallelism (SP) within the LLM backbone unit and automatically splits sequences within its SP group.
+Furthermore, DistTrain supports expert parallelism (EP) for the LLM backbone.
+Since EP and TP both perform parallel computation and communication within one layer, our subsequent formulation involving TP remains valid when TP is replaced with EP.
+The detailed implementation of parallelism unit is described in §6.
+4.2. Problem Formulation
+With disaggregated model orchestration, we are able to adaptively orchestrate the three modules.
 The problem now lies in determining the optimal resource allocation and parallelism strategy to minimize the training time per iteration.
-Exhaust search is infeasible due to the large search space, particularly in large cluster.
+Exhaust search is infeasible due to the combinatorial search space, particularly for large clusters.
 One strawman solution is to allocate the resources proportional to the model flops of each module.
 However, this method falls short as it overlooks complex patterns in parallelism training.
-Before diving into disaggregated model orchestration, we first formulate the optimization problem.
+Different DP and TP sizes in modality encoder could lead to different computation times while the model flops are the same.
+Before diving into our solution, we first formulate the optimization problem.
 LLM backbone.
-We begin by formulating the LLM backbone, focusing on the forward pass as the backward pass time mirrors this.
-In LLM training, microbatch size is set to one to prevent GPU memory overflow.
+We begin by formulating the LLM backbone, focusing on the forward pass as the backward pass mirrors it.
+In production LLM training, the microbatch size is set to a small number to prevent GPU memory overflow.
+DistTrain sets the microbatch size to a predefined constant, .
 Assume the global batch size for one iteration as and the TP size of the LLM backbone as .
-Let the PP and DP size of LLM backbone be and .
-The number of GPUs allocated to LLM backbone is .
-Let the forward time (including communication time) of the entire LLM be , where represents forward time function.
-Therefore, the forward time of one PP stage for one microbatch is .
+Let the PP and DP size of the LLM backbone be and .
+The LLM is divided into equal PP stages, where each PP stage consists of some homogeneous transformer layers.
+The number of GPUs allocated to the LLM backbone is .
+Let the forward time (including communication time) of the entire LLM for one sample be , where represents the forward time function.
+Therefore, the forward time of one PP stage for one microbatch with samples is since PP partitions LLM backbone’s homogeneous layers sequentially.
 Besides, the number of microbatches per iteration is .
@@
 Let the TP size be and the DP size be .
-The number of GPUs allocated to modality encoder is .
+The number of GPUs allocated to the modality encoder is .
 The microbatch size is which is determined by the LLM backbone.
-Let the forward time (including communication time) of the entire modality encoder be .
+Let the forward time of the entire modality encoder for one training sample be .
 The forward time of one PP stage for one microbatch in the modality encoder is .
@@
 Based on the preceding analysis, we next define the objective function for the optimization problem, i.e., the training time of one iteration.
-As shown in Figure 10, we demonstrate the pipeline of forward pass in multimodal LLM training.
+Figure 10 shows the pipeline of forward pass in multimodal LLM training.
 The LLM backbone comprises two PP stages, whereas the modality encoder and generator each consist of one PP stage.
-The training process is categorized into two phases: warm-up phase and steady phase.
+The forward pass is categorized into two phases: warm-up phase and steady phase.
 The warm-up phase spans from the initiation to the completion of the first microbatch to populate the pipeline.
-This phase’s duration is calculated as , which is formulated as follows.
-The steady phase’s duration is dominated by the maximal computation time among PP stages, which is calculated as , where is the number of microbatches per iteration. It is formulated as:
+The duration of this phase is calculated as , which is formulated as follows.
+The duration of steady phase is dominated by the maximal computation time among PP stages, which is calculated as , where is the number of microbatches per iteration. It is formulated as:
 Therefore, the objective function is to minimize .
-For the backward pass, the objective function remains analogous to that of the forward pass.
+Considering backward pass with the similar computation pattern, the objective function is similar to that of the forward pass.
 Adjustments are made by changing , and from forward time functions to the sum functions of forward and backward time.
-This formulation holds for GPipe and 1F1B.
-We will retrofit the formulation to adapt to VPP later.
+This formulation holds for GPipe (Huang et al., 2019) and 1F1B (Fan et al., 2021).
+We will retrofit the formulation to adapt to VPP (Narayanan et al., 2021) later.
 TP communication is incorporated into the functions , and , which are calibrated through interpolation from actual trials.
 The communication time of DP and PP is modeled as the communication volume divided by the bandwidth.
-If the DP sizes are fixed, the DP communication time remains constant.
-For PP, the communication time equals the PP size multiplied by the communication time of a single layer’s output tensor.
 Constraints.
@@
 The first constraint is the resource constraint.
-The number of GPUs allocated to each module should be where is the total number of GPUs in the cluster.
-The second constraint involves memory.
-We consider LLM backbone.
+The number of GPUs allocated to each module should be where is the total number of GPUs in the cluster and are the number of GPUs allocated to each module in multimodal LLM.
+The second constraint involves GPU memory.
+We first consider LLM backbone.
 Memory allocation involves four parts: model parameters, gradients, optimizer states, and activation states.
-The memory of model parameters and gradients on one GPU is calculated as: , where denotes total memory for the LLM parameters and gradients.
-The memory for optimizer states on one GPU (with ZeRO-1 optimization) is: , where denotes the total memory for the optimizer states.
+The memory of model parameters and gradients on one GPU is calculated as: , where denotes the total memory for the LLM parameters and gradients.
+The memory for optimizer states on one GPU (with ZeRO-1 optimization (Rajbhandari et al., 2020)) is: , where denotes the total memory for the optimizer states.
 ZeRO-1 partitions the optimizer states across DP groups.
-The peak memory for activation states on one GPU is: , with representing the memory needed for one microbatch of activation states across the entire LLM.
+The peak memory for activation states on one GPU is: , with representing the memory needed for one microbatch (with microbatch size ) of activation states across the entire LLM.
 In 1F1B, the first PP stage requires storage for microbatches of activation states.
-We eschew using GPipe in DistTrain since GPipe consumes more memory.
-The memory constraint ensures the sum of the four memory parts on one GPU doesn’t exceed GPU capacity.
-As for modality encoder and generator, the formulation is similar.
-5.2 Disaggregated Model Orchestration
-The optimization problem is non-convex, with , and the DP, TP sizes as positive variables.
-is predefined by the user.
-Solving this with an exhaust search algorithm to explore all possible variable values is impractical due to the extensive search space, particularly in large clusters.
-Designing an efficient algorithm that quickly identifies the optimal resource allocation and parallelism strategy is a significant challenge.
+We do not use GPipe in DistTrain since it consumes more memory without offering better training efficiency compared to 1F1B.
+The memory constraint ensures the sum of the four memory parts on one GPU does not exceed GPU capacity.
+As for the modality encoder and generator, the formulation is similar.
+4.3. Adaptive Model Orchestration
+The problem is non-convex, with , and the DP, TP sizes as positive variables.
+Solving this with an exhaust search is impractical due to the large search space, particularly in large clusters.
+Designing an efficient algorithm that adaptively identifies the optimal resource allocation and parallelism strategy based on the training task poses a hard problem.
 Convex optimization.
@@
 We confine the TP size to on an NVIDIA GPU node with 8 GPUs and adjust the DP size as a factor of to balance the computation across DP groups.
+If expert parallelism (EP) is used instead of TP, we confine EP size to factors of cluster size.
 The PP size of the LLM backbone is calculated as .
 The set of possible parallelism strategies is a manageable and finite set, i.e., the Cartesian product of TP and DP size.
-This allows us to enumerate all feasible TP and DP sizes and transform the original optimization problem into a set of simplified problems.
-In the simplified problem, the objective function is maximal and additional of the functions: , and , where are positive.
+This allows us to enumerate all feasible TP and DP sizes, and transform the original optimization problem into a set of simplified problems.
+In the simplified problem, the objective function is maximal and additional of the functions: , and , where are positive variables.
 Therefore, the objective function is convex.
 Similarly, the constraint functions are also convex.
-As a result, the simplified optimization problem is convex and can be efficiently solved to optimality by existing solvers [34, 35].
-The algorithm, detailed in Algorithm 1, efficiently finds the optimal resource allocation and parallelism strategy.
-Virtual pipeline parallelism (i.e., interleaved 1F1B) reduces the warm-up time by dividing model into finer-grained virtual PP (VPP) stages.
+As a result, the simplified optimization problem is convex and can be efficiently solved to optimality by off-the-shelf solvers (Grant and Boyd, 2014; Virtanen et al., 2020).
+The algorithm adaptively determines the optimal resource allocation and parallelism strategy.
+Virtual pipeline parallelism (Narayanan et al., 2021) (i.e., interleaved 1F1B) reduces the warm-up time by dividing model into finer-grained virtual PP (VPP) stages.
 Each PP stage contains VPP-size VPP stages.
@@
 To align our formulation with VPP, we proportionally reduce the warm-up time based on VPP-size.
-6 Addressing Data Heterogeneity
-Disaggregated training enables preprocessing the training data with negligible runtime overhead.
-Based on this, we introduce disaggregated data reordering, seamlessly integrated into data preprocessing, to address data heterogeneity without additional overhead.
-Initially, we present intra-microbatch reordering to eliminate stragglers across DP groups.
-We then introduce inter-microbatch reordering to minimize pipeline bubbles caused by the straggler microbatches.
-The combination of the two reordering, i.e., disaggregated data reordering, effectively addresses the data heterogeneity.
-We emphasize that these two reordering algorithms are integrated into the disaggregated data preprocessing.
-This ensures that the complex reordering does not interfere with the GPU training process.
-6.1 Intra-microbatch Reordering
+5. Addressing Data Heterogeneity
+To address data heterogeneity, we first introduce disaggregated data preprocessing that separates data preprocessing from training.
+This eliminates the interference between data processing and training and enables negligible overhead of data processing.
+We then introduce disaggregated data reordering to address data heterogeneity.
+We use intra-microbatch reordering to eliminate stragglers across DP groups, and inter-microbatch reordering to minimize pipeline bubbles and hide the training stragglers into the training pipeline as much as possible.
+5.1. Disaggregated Data Preprocessing
+As discussed in §2.3, preprocessing high-volume multimodal data (e.g., high resolution images) incurs substantial overhead.
+DistTrain disaggregates data preprocessing from training with a producer-consumer model.
+The producer, operating on dedicated CPU nodes, fetches data from the distributed file system and preprocesses training data asynchronously.
+The consumer, i.e., the main training process on dedicated GPU nodes, receives the preprocessed data for training.
+The producer and consumer communicate through RPC calls, and use RDMA for lower latency if available.
+The disaggregation enables flexible and elastic data preprocessing and provides opportunities for data reordering without incurring additional overhead.
+Leveraging this opportunity, we incorporate data reordering into data preprocessing to mitigate training stragglers caused by data heterogeneity.
+The reordering includes two levels: intra-microbatch and inter-microbatch.
+We emphasize that disaggregated data preprocessing ensures that the complex reordering does not interfere with the GPU training or impose extra overhead.
+5.2. Intra-microbatch Reordering
 Insight.
-To address the first subproblem, intra-microbatch stragglers, we identify the straggler by pinpointing the DP group with the largest training samples.
-As illustrated in Figure 6, the first DP group becomes a straggler as it contains the two largest training samples.
-To neutralize this imbalance, we propose reordering the training samples within the global batch by size.
-Specifically, as depicted in Figure 11, we reorder the training samples into the sequence .
-This strategy effectively distributes the computational load more evenly and improves the scalability.
-It eliminates the straggler without breaking the synchronous training semantics [31], as the reordering occurs within an individual global batch.
+To address intra-microbatch stragglers, we first identify the straggler by pinpointing the DP group with the largest training samples.
+As shown in Figure 6, the first DP group becomes a straggler as it contains the two largest training samples.
+Therefore, we reorder the training samples within the global batch by size.
+Specifically, as depicted in Figure 11, we reorder the training samples into the sequence , which distributes the computation more evenly.
 Intra-microbatch Reordering.
-Leveraging this insight, we propose intra-microbatch reordering to balance computational load and improve the overall scalability.
-Formally, the challenge involves minimizing the maximum computation time among DP groups.
-This problem corresponds to the NP-hard problem, multiway number partitioning [36], which aims to minimize the largest sum.
-There is no known polynomial-time algorithm capable of determining the optimal solution.
-Given the substantial batch size in production training, the algorithm employed must be lightweight and efficient.
-Consequently, we adopt the greedy number partitioning algorithm, which guarantees an approximation ratio of [37].
+Leveraging this insight, we propose intra-microbatch reordering to balance the computational load within a global batch.
+Formally, the goal is to minimize the maximum computation time across data parallelism (DP) groups.
+This problem maps to the multiway number partitioning problem (Korf, 2009), which is NP-hard.
+Given the problem’s complexity and the large batch sizes common in production, finding an optimal solution is computationally infeasible.
+We therefore employ a lightweight, greedy partitioning algorithm that yields an approximation ratio of less than 4/3 (Barman and Krishnamurthy, 2020).
+Importantly, this reordering only occurs within the global batch, which only affects the sequence of gradient accumulation.
+Since gradient accumulation (i.e., summation) is a commutative operation, the intra-microbatch reordering preserves the training’s convergence semantics.
 The detailed algorithm is summarized in Algorithm 1.
 The function receives the original training samples and DP size .
-This algorithm first sorts the training samples in ascending order by the sample size (line 3).
+This algorithm first sorts the training samples in descending order by the sample size (line 3).
 Then, it loops over the training samples and assigns the sample to the DP group with the current lowest computational load (line 6-8).
@@
 The algorithm has a time complexity of .
-6.2 Inter-microbatch Reordering
-The second subproblem is the inter-microbatch straggler.
+5.3. Inter-microbatch Reordering
 As we discussed in §2.3, data heterogeneity leads to varied computation times across microbatches within the modality encoder and generator.
 The straggler microbatch prolongs the training by creating large pipeline bubbles.
-In the context of 1F1B pipeline scheme, the overall iteration time is primarily governed by pipeline bubbles and the computation time at the first PP stage of the modality encoder, as illustrated in Figure 12.
+In the context of 1F1B pipeline scheme, the overall iteration time is primarily decided by pipeline bubbles and the computation time at the first PP stage of the modality encoder, as illustrated in Figure 12.
 Let the PP size be and the number of microbatches be ( and in Figure 12).
 Typically, is larger than to reduce the proportion of time spent in the warm-up and cool-down phases.
-We abstract an innovative concept, pipeline intervals, at the first PP stage.
-As shown in Figure 12, these intervals are typically filled with forward pass, except for the last intervals (i.e., , , and ).
+To help describe our solution, we first introduce a concept called pipeline intervals at the first PP stage.
+As shown in Figure 12, these intervals are typically filled with the forward pass, except for the last intervals (i.e., , , and ).
 Straggler microbatches in either the encoder or generator prolong these intervals or increase the unfilled area (i.e., bubble).
+Designing an efficient runtime algorithm to minimize pipeline bubbles presents a significant challenge.
 Insights.
@@
 As shown in Figure 12, the last intervals (i.e., to ) remain unfilled.
-These intervals become the pipeline bubbles and prolong the training iteration.
+These intervals become the pipeline bubbles and increase the training iteration.
 We observe a positive correlation between the volume of and the size of the microbatch.
 The size refers to the computation time of the microbatch in modality encoder and generator.
-For instance, is significantly larger than and since the microbatch is the largest.
-By strategically reordering the training samples to position the smallest microbatches at the end, we are able to reduce these unfilled intervals (i.e., pipeline bubbles).
-The second insight involves minimizing the unfilled area of left intervals.
+For instance, is significantly larger than and since the microbatch (i.e., d) is the largest.
+By strategically reordering the training samples to position the smallest microbatches at the end, we are able to reduce these unfilled intervals’ volume.
+The second insight involves minimizing the unfilled area of the left intervals, which hides the training stragglers into the training pipeline as much as possible.
 The left intervals (i.e., to ) are filled with the forward pass.
@@
 For subsequent intervals, is filled by the forward pass.
-By evaluating the volume of , we place the microbatches, whose forward time most closely matches this volume, at the corresponding position, to minimize the unfilled area (i.e., pipeline bubbles).
+However, the forward pass may not perfectly match the interval volumes, leading to unfilled areas.
+By evaluating the volume of , we place the microbatches, whose forward time most closely matches this volume, at the corresponding position, to minimize the unfilled area.
 Inter-microbatch Reordering.
-Based on the two insights, we design runtime inter-microbatch reordering to minimize the pipeline bubbles.
-This algorithm is designed for 1F1B pipeline scheme.
+Based on the preceding insights, we propose a runtime algorithm for inter-microbatch reordering designed to minimize pipeline bubbles.
+This algorithm is specifically developed for the 1F1B pipeline schedule.
 We will retrofit the algorithm to VPP (i.e., interleaved 1F1B) later.
 Algorithm 2 summarizes the pseudo code.
-The function receives the original order of microbatches and the PP size .
-It stores reordered microbatches in and pending microbatches in (line 2).
-Early in the process, the smallest microbatch is placed at first to activate all pipeline stages promptly (line 3).
-Subsequently, it selects the smallest microbatches and places them at the end to minimize unfilled intervals (line 4 and line 12).
-It then loops over the remaining microbatches (line 5-11).
-In each loop iteration, it calculates the volume of the interval through the function .
-For the first interval, it selects microbatches that closely match the interval volume in sum forward time; for others, it selects a single microbatch whose forward time aligns closely with the interval volume.
+The function accepts the initial microbatch sequence and the PP size .
+The reordering process begins by scheduling the smallest microbatch first to ensure all pipeline stages are activated promptly (line 3).
+Subsequently, it reserves the smallest remaining microbatches for the end of the sequence to minimize unfilled intervals (lines 4 and line 12).
+The main loop (lines 5-11) then iterates through the remaining microbatches to fill the pipeline bubbles.
+In each loop iteration, it computes the interval size via the function and applies a heuristic to select microbatches that best fit this size.
+For the first interval, it greedily selects microbatches whose aggregate forward time most closely matches the interval size; for all subsequent bubbles, it selects the single best-fitting microbatch.
 This loop ensures maximal filling of the remaining intervals, which minimizes pipeline bubbles.
+Crucially, this reordering only alters the data sequence within the local batch of a single DP rank per training iteration.
+Therefore, analogous to intra-microbatch reordering, it only affects the sequence of gradient accumulation and consequently preserves the training’s convergence semantics.
 The functions and operate with a time complexity of .
-The function calculates interval volumes using the current order .
+The function calculates interval size using the current order .
 This calculation is facilitated by a dynamic programming algorithm that utilizes a recursive formula derived from pipeline dependencies.
-Specifically, each microbatch’s start time depends on two factors: the completion of the preceding microbatch on the same device and the availability of input data from the upstream microbatch.
+Specifically, the start time of each microbatch depends on two factors: the completion of the preceding microbatch in the same pipeline stage and the availability of input data from the upstream microbatch.
 Consequently, the end time of each microbatch is determined by the maximum of these two dependencies plus its own computation time.
@@
 Virtual pipeline parallelism (i.e., interleaved 1F1B) also follows the one forward and one backward pipeline scheme to reduce the memory footprint.
-The fundamental insights of our algorithm apply universally to any 1F1B-based pipeline, including VPP.
+The fundamental insights of our algorithm apply to any 1F1B-based pipeline, including VPP.
 We adapt the algorithm by computing multiple (i.e., VPP size) intervals and filling them with the corresponding number of forward passes from a single microbatch.
-7 System Implementation and Optimization
-We implement DistTrain with 6.3K lines of code in Python and C++, and integrate it with Megatron-LM [17], a state-of-the-art training framework for large language models.
-DistTrain is able to support multimodal LLM training with a variety of modalities on a large-scale GPU cluster (e.g., thousands of GPUs used in §8).
-DistTrain leverages a distributed file system to store the training data and model checkpoints.
-It handles failures by automatically recovering the training process from the latest checkpoint.
-DistTrain manager.
-DistTrain’s training manager, implemented as a Python script on a dedicated CPU node, formulates the disaggregated model orchestration problem using Disciplined Convex Programming [38].
-It employs the CVX solver [39] to efficiently solve this problem within milliseconds.
-The manager then records the optimal resource allocation and parallelism strategy to a configuration file, which the Kubernetes controller uses to launch the training task.
-DistTrain initializer.
-DistTrain incorporates parallelism unit that manages specific parallelism strategies for each module in multimodal LLMs.
-Parallelism unit is initialized using PyTorch Distributed [40] library with NCCL as the communication backend, except for the TP communication where we use StepCCL instead.
-DistTrain shards the model parameters in accordance with the established parallelism strategy.
+6. System Implementation
+We implement DistTrain with 6.3K lines of code in C++ and Python, and integrate it with Megatron-LM (Shoeybi et al., 2019).
+DistTrain handles failures by automatically recovering the training from the latest model checkpoint.
+To mitigate Tensor Parallelism (TP) overhead, we developed StepCCL, an in-house collective communication library.
+By leveraging the DMA engine for data transfers, StepCCL overlaps communication with computation, which frees the core Streaming Multiprocessors (SMs) to execute computation kernels (e.g., GEMM) without performance interference.
+A detailed description and evaluation are in §A.1.
+DistTrain manager and initializer.
+DistTrain’s training manager, running on a dedicated CPU node, formulates the disaggregated model orchestration problem using Disciplined Convex Programming (Grant et al., 2006).
+It employs the CVX solver (cvx, 2024) to solve this problem within a second.
+The manager records the optimal resource allocation and parallelism strategy to a configuration file, which the Kubernetes controller uses to launch the training task.
+As for the initializer, DistTrain uses PyTorch Distributed (tor, 2024) library to initialize the communication groups.
+The initialization of each module in a multimodal LLM is implemented through parallelism units.
 Parallelism unit.
-As we discussed in §4.1, the GPU training disaggregation is implemented through a specific module: parallelism unit.
-When initializing the distributed training, DistTrain first establishes the communication groups within one parallelism unit.
-Each GPU process has a global rank and a local rank within the unit, which facilitates the distributed initialization.
-Then, it initializes the communication broker to establish the PP communication between adjacent parallelism units.
-All communication traffic between parallelism units is routed via the communication broker.
-We implement the communication broker by modifying the batched send and receive operations in Megatron-LM to separate operations.
-This allows flexible communication between multiple upstream and downstream GPU processes.
-The communication broker adjusts the communication data by concentrating and scattering the data as needed while maintaining the data order.
-Strategically located on the GPU of the last PP stage in the upstream unit or the first PP stage in the downstream unit, the communication broker avoids additional communication overhead.
-Besides, the number of communication brokers between two units is determined by the greatest common divisor of the two parallelism units’ DP sizes to maximize the communication bandwidth.
-Moreover, in the Megatron-LM, the reliance on synchronous communication compels upstream stages to pause until downstream stages fully receive the data.
-This introduces unnecessary dependencies in the pipeline.
-To alleviate this, we implement asynchronous send operations that eliminate these superfluous dependencies and redesign the communication topology to prevent potential deadlocks.
-Mitigating TP overhead with StepCCL.
-Tensor Parallelism (TP) is commonly adopted to facilitate training large Transformer-based models with multiple GPUs connected with high bandwidth (e.g., with NVLinks).
+As discussed in §4.1, disaggregated model orchestration is implemented via a dedicated module: parallelism unit.
+During distributed training initialization, DistTrain first establishes communication groups within a parallelism unit.
+Each GPU process possesses a global and a local rank within its unit, facilitating distributed initialization.
+Subsequently, DistTrain initializes a communication broker to establish PP communication between adjacent parallelism units.
+All inter-unit communication traffic is routed via the communication broker.
+The communication broker is implemented by modifying Megatron-LM’s batched send/receive operations into discrete operations.
+It enables flexible communication between multiple upstream and downstream GPU processes by concentrating and scattering data as needed, while preserving data order.
+The communication broker employs a decentralized design, residing on the GPU of either the last PP stage in an upstream unit or the first PP stage in a downstream unit.
+To maximize communication bandwidth, the number of brokers between two units is determined by the greatest common divisor of their respective DP sizes.
+Consequently, the total inter-unit bandwidth scales effectively with the training workload, preventing the communication broker from becoming a training bottleneck.
+Moreover, Megatron-LM’s reliance on synchronous communication compels upstream stages to pause until downstream stages fully receive data, introducing unnecessary pipeline dependencies.
+To mitigate this, we implement asynchronous send operations, eliminating these superfluous dependencies, and redesign the communication topology to prevent potential deadlocks.
+7. Evaluation
+In this section, we first use large-scale experiments to evaluate the overall performance improvements of DistTrain over Megatron-LM.
+Next, we use an ablation study to deep dive into DistTrain and show the effectiveness of each technique.
+Finally, we provide a case study to further evaluate DistTrain.
+Setup.
+Our experiments are conducted on a production cluster, with each node equipped with 8 NVIDIA Ampere GPUs.
+GPUs within one node are interconnected by 300GB/s (bidirectional) NVLink, while nodes are connected by 4*200 Gbps RDMA network based on RoCEv2 with rail-optimized topology.
+The overall experiments use up to 1296 GPUs, and the ablation study utilizes up to 96 GPUs.
+We use PyTorch 2.1.2 and NVIDIA CUDA 12.2.
+As for the modality, we focus on images and texts.
+We emphasize that DistTrain is also compatible with other modalities.
+Our evaluation focuses on images (i.e., vision).
+This choice is motivated by the prevalence of vision as a common non-textual modality in multimodal LLMs and the increasing prominence of capable vision-based MLLMs (e.g., Gemini (gem, 2024) and Qwen2.5-VL (Bai et al., 2025)).
+Models.
+For the LLM backbone, we choose the representative LLM architecture, Llama3 (lla, 2024), which is widely used in both academia and industry.
+Table 2 lists the detailed model configurations.
+For encoder and generator, we use ViT-Huge (vit, 2024) (0.63B) and SD 2.1 (sd, 2024) (1B) respectively.
+The encoder and generator in our experiments are representative of state-of-the-art MLLM.
+The encoder (i.e., ViT-Huge) in our experiment aligns with the encoders of Qwen2.5-VL and Seed1.5-VL in terms of both model architecture and size.
+The generator (i.e., SD) is also adopted in many MLLMs (e.g., Step1X-Edit (Liu et al., 2025) and EMU (Sun et al., 2023b)) with generation capabilities.
+The three LLM backbones (i.e., Llama3-7B, Llama3-13B, and Llama3-70B) are paired with ViT-Huge and SD to form multimodal LLMs denoted as MLLM-9B, MLLM-15B, and MLLM-72B.
+For the large multimodal LLM (i.e., MLLM-72B), we use high image resolution (i.e., 10241024) for generation since the large LLM is able to process more context information.
+For small models, we use low image resolution (i.e., 512512).
+Models | # of | Hidden | FFN Hidden | # of | # of |
+Layers | Size | Size | Heads | Groups | |
+Llama3-7B | 32 | 4096 | 11008 | 32 | 32 |
+Llama3-13B | 40 | 5120 | 13824 | 40 | 40 |
+Llama3-70B | 80 | 8192 | 28672 | 64 | 8 |
+Datasets.
+For our experiments, we use the representative open-source dataset, LAION-400M (Schuhmann et al., 2021).
+We generate training data by interleaving the image and text subsequences, forming input sequences up to 8192 tokens long.
+This dataset is also employed in our production multimodal LLM training.
+As detailed in §2.3, each training sample includes a varying number of image tokens and text tokens, which introduces data heterogeneity in multimodal LLM training.
+Metrics.
+We use Model FLOPs Utilization (MFU) as the primary metric.
+MFU measures the percentage of GPU FLOPs that are effectively utilized during training.
+We also use the training throughput to evaluate the training speed of DistTrain.
+Since DistTrain and Megatron-LM may utilize different numbers of GPUs due to varying model orchestration strategies, we also indicate the number of GPUs used in each experiment.
+Baselines.
+We focus on training multimodal LLMs rather than traditional multimodal models like CLIP or LiT.
+DistMM (Huang et al., 2024) efficiently handles traditional multimodal training but lacks support for pipelines involving encoders, LLM backbones, and generators in multimodal LLMs.
+Multimodal LLMs combine multimodal models and LLMs to understand and generate high-quality multimodal data auto-regressively.
+Megatron-LM is selected as the primary baseline due to its established role as a highly optimized, open-source framework for LLM training that also provides supports for multimodal models.
+For ablation study of model orchestration, we integrate DistMM’s model orchestration strategy (resource allocation by model size and FLOPs) into DistTrain, naming this baseline DistMM∗.
+Note that DistMM does not support multimodal LLM training natively.
+DistMM∗ only uses its orchestration strategy, with all other techniques from DistTrain.
+7.1. Overall Performance
+We first compare the overall performance of DistTrain against Megatron-LM on a large GPU cluster (up to 1296 GPUs).
+We retrofit Megatron-LM to support multimodal LLM training by integrating modality encoder and generator into the training pipeline.
+Megatron-LM employs monolithic model orchestration as described in §2.1.
+In Megatron-LM, we set the PP size of the LLM backbone to 1, 2, and 10 for Llama3-7B, Llama3-13B, and Llama3-70B, respectively.
+PP size is set to 1 for modality encoder and generator.
+We set TP size to 8, as each node consists of 8 GPUs connected by high-bandwidth NVLink.
+As for DistTrain, the parallelism strategy is determined by disaggregated model orchestration.
+In our experiments, one GPU is able to facilitate training ViT and SD.
+We replicate the modality encoder and generator across the GPUs within the TP group to process different images, whereas TP itself is not used.
+We set the global batch size to 1920.
+The experimental results are shown in Figure 13 and Figure 14.
+Figure 13 shows the MFU.
+Figure 14 shows the training throughput.
+Due to the different model orchestration strategies of DistTrain and Megatron-LM, DistTrain uses 1056, 1216, and 1176 GPUs for MLLM-9B, MLLM-15B, and MLLM-72B, respectively.
+Megatron-LM uses 1296, 1280, and 1152 GPUs, respectively.
+Notably, within a total budget of 1296 GPUs, DistTrain intentionally allocates fewer resources in some cases because adding more GPUs yields no further improvements in training throughput.
+This resource efficiency frees the remaining GPUs for concurrent tasks such as fine-tuning or inference.
+We summarize the findings as follows.
+As shown in Figure 13, DistTrain achieves 51.8%-54.7% MFU in large-scale multimodal LLM training.
+This performance closely approximates that of state-of-the-art unimodal (i.e., text) LLM training (Jiang et al., 2024), which demonstrates the effectiveness of DistTrain in addressing the model and data heterogeneity in multimodal LLM training.
+DistTrain significantly outperforms Megatron-LM, delivering 1.7–2.8 the MFU and 1.7–2.2 the training throughput when training MLLM-9B and MLLM-15B with a similar number of GPUs.
+These performance gains largely stem from DistTrain’s disaggregated model orchestration.
+Megatron-LM’s monolithic strategy often leads to GPU underutilization, since it assigns too many GPUs to the modality encoder and generator.
+In contrast, DistTrain adaptively adjusts model orchestration based on specific model and data demands.
+Additionally, DistTrain’s disaggregated data preprocessing technique further improves efficiency.
+In the MLLM-72B training scenario, DistTrain outperforms Megatron-LM by 1.2 on MFU and 1.3 on training throughput with a similar number of GPUs.
+The high image resolution increases the execution time of the multimodal module, which introduces pipeline bubbles in LLM backbone.
+DistTrain addresses this by allocating additional GPUs to these modules to balance the pipeline.
+The disaggregated data preprocessing strategy continues to mitigate data heterogeneity, thereby increasing training efficiency.
+7.2. Ablation Study
+In this subsection, we perform microbenchmarks to evaluate the effectiveness of each DistTrain’s technique and utilize up to 96 GPUs.
+We reduce the cluster scale due to our limited budget.
+We set the global batch size of training to 128, 64, and 40 for MLLM-9B, MLLM-15B, and MLLM-72B, respectively.
+Disaggregated model orchestration.
+We evaluate the MFU and training throughput using the model orchestration strategies of DistTrain, Megatron-LM, and DistMM∗.
+DistMM∗ allocates GPUs according to the computation demands (flops) of each module.
+DistTrain uses 96, 80, and 82 GPUs for MLLM-9B, MLLM-15B, and MLLM-72B, respectively.
+Megatron-LM uses 96, 96, and 96 GPUs, while DistMM∗ uses 88, 76, and 90 GPUs.
+The experimental results are shown in Figure 15.
+DistTrain consistently outperforms the baseline strategies, achieving 1.3–2.7 higher MFU and 1.4–2.7 higher training throughput.
+Although DistMM∗ outperforms Megatron-LM, it still lags behind DistTrain since it neglects the intricate performance model (§4.2) of multimodal LLM training.
+DistTrain’s disaggregated model orchestration optimally balances computational loads across the three modules and achieves high resource utilization.
+In production training, the running time of disaggregated model orchestration is less than one second, negligible compared to the days required for full training.
+(a) MFU. (b) Throughput.
+We also evaluate the running time of DistTrain’s disaggregated model orchestration to determine the optimal resource and parallelism configurations under different training settings, as detailed in Table 3.
+The algorithm completes in under one second.
+The overhead is negligible compared to the days or even weeks required for overall training.
+Disaggregated data preprocessing.
+We evaluate the effectiveness of DistTrain’s disaggregated data preprocessing by comparing it against Megatron-LM’s data preprocessing, while keeping other components the same.
+Megatron-LM’s data preprocessing uses random ordering for training data.
+The effectiveness is gauged through metrics such as MFU and training throughput.
+We use the optimal resource allocation and parallelism strategy decided by DistTrain’s disaggregated model orchestration.
+Given that the model orchestration strategy remains unchanged, we do not indicate the number of GPUs.
+The experimental settings are the same as those in the experiment to evaluate disaggregated model orchestration.
+The results are shown in Figure 16.
+DistTrain consistently outperforms the baseline, achieving 1.03-1.11 higher MFU and throughput.
+The performance gap becomes more pronounced as the model size decreases.
+This is because the smaller model size leads to a higher data parallelism (DP) size, which causes more intra-microbatch heterogeneity.
+In essence, DistTrain’s two-level data reordering effectively mitigates data heterogeneity and enhances training efficiency.
+We do not measure the running time of the reordering algorithm as it operates on dedicated CPU nodes asynchronously.
+The data preprocessing overhead with and without disaggregation is evaluated in §7.3.
+Model | # of | Global | Algorithm |
+GPUs | Batch Size | Overhead | |
+MLLM-72B | 1296 | 1920 | 922ms |
+MLLM-72B | 648 | 960 | 641ms |
+MLLM-72B | 324 | 480 | 441ms |
+MLLM-72B | 112 | 240 | 133ms |
+7.3. Case Study
+In this subsection, we first evaluate the overhead of data preprocessing. We then evaluate DistTrain under different frozen training settings.
+(a) MFU. (b) Throughput.
+Overhead of data preprocessing.
+We conduct an experiment to evaluate the overhead of data preprocessing, including decompression and reordering.
+Setting the DP size to one, we measure the average data preprocessing time per iteration on the GPU training side.
+We then compare data preprocessing time with and without disaggregated data preprocessing and use varying numbers of images and different image resolutions for one training iteration.
+The results, depicted in Figure 17, indicate disaggregating data preprocessing from training reduces preprocessing time from seconds to milliseconds.
+The first parameter in the x-axis represents the number of images, while the second parameter denotes the image resolution.
+In production training (§7.1), iteration times range from seconds to tens of seconds.
+Without disaggregation, preprocessing overhead is in seconds, significantly interferes with training.
+With disaggregated data preprocessing, the overhead reduces to milliseconds, which is negligible relative to total iteration time.
+(a) All modules freezing. (b) Encoder-only training. (c) LLM-only training. (d) Generator-only training.
+(a) All modules freezing. (b) Encoder-only training. (c) LLM-only training. (d) Generator-only training.
+Frozen training.
+In real production training, some modules in multimodal LLM are frozen to stabilize training loss and enhance model effectiveness during different training phases (Yin et al., 2023).
+We conduct a frozen training experiment under four specific training settings: complete module freezing (i.e., training projectors only), encoder-only training, LLM-only training, and generator-only training.
+In these scenarios, frozen modules neither compute weight gradients in backward pass nor update weights.
+The frozen modules still perform computations in forward pass.
+All other experimental setup aligns with those detailed in §7.2.
+DistTrain uses 96, 80, and 82 GPUs for MLLM-9B, MLLM-15B, and MLLM-72B under all three frozen settings.
+Megatron-LM uses 96, 96, and 96 GPUs for MLLM-9B, MLLM-15B, and MLLM-72B under all three frozen settings.
+The experimental results are in Figure 18 and Figure 19.
+DistTrain consistently outperforms Megatron-LM across all frozen training configurations, achieving 1.4–2.9 higher MFU and 1.2–2.9 higher training throughput.
+This pronounced performance gap underscores the challenges posed by Megatron-LM’s monolithic model orchestration in more complex training environments.
+In contrast, DistTrain adaptively adjusts model orchestration based on training settings and achieves high resource utilization.
+8. Discussion
+Parallelism strategies.
+Many studies (Zheng et al., 2022; Wang et al., 2019; Jia et al., 2019) optimize parallelism strategies to accelerate deep learning model training.
+However, they fall short of holistically considering data, pipeline, and tensor parallelism in multimodal LLM training, due to the large search spaces in large clusters.
+This limitation results in inefficient and sub-optimal parallelism strategy.
+Additionally, these methods generally assume homogeneity of training data—all training samples consume the same amount of computation.
+This assumption does not hold for multimodal LLMs where training samples of different modalities vary significantly in volume and computation demand.
+DistTrain leverages the specific training pattern of multimodal LLMs, and formulates a model orchestration problem for multimodal LLMs that integrates tensor, pipeline, and data parallelism simultaneously.
+Besides, DistTrain leverages disaggregated data reordering to address data heterogeneity.
+Sequence and expert parallelism.
+Sequence parallelism (SP) (Li et al., 2021) is designed to partition the training sequence into multiple subsequences for parallel training.
+It addresses the challenges of processing long sequences in LLMs.
+Expert parallelism (EP) (Liu et al., 2023b), specifically devised for mixture-of-experts (MoE) LLMs (Du et al., 2022), enables parallel training of multiple feedforward network (FFN) experts.
+These parallelism strategies are orthogonal to multimodal LLM training.
+In DistTrain, both SP and EP are integrated into the LLM backbone training.
+Heterogeneous hardware.
+By disaggregating three modules (i.e., the modality encoder, LLM backbone, and generator) for multimodal LLM training, DistTrain supports using heterogeneous hardware for different modules to achieve varius goals like improving performance, reducing cost and increasing energy efficiency.
+This strategy assigns distinct computational workloads to the most appropriate hardware.
+For instance, ViT encoder is significantly less compute-intensive than LLM backbone.
+Consequently, we can place ViT encoder on more economical GPUs (e.g., NVIDIA L20).
+Such benefits are demonstrated in other LLM systems such as MegaScale-Infer (Zhu et al., 2025) and Splitwise (Patel et al., 2024).
+Disaggregated data preprocessing.
+The principal technique in DistTrain’s disaggregated data preprocessing is data reordering that addresses training stragglers caused by data heterogeneity.
+Although data reordering does not inherently necessitate disaggregation, we adopt disaggregation to mitigate the preprocessing overhead and improve system elasticity (e.g., allocating any arbitrary number of CPUs for preprocessing tasks).
+This approach is analogous to resource disaggregation in other systems scenarios, which aims to improve system elasticity and efficiency (Shan et al., 2018; Wang et al., 2022a).
+9. Related Work
+LLM training.
+Many efforts have been made to optimize LLM training.
+For LLM pre-training, Megatron-LM (Shoeybi et al., 2019) and DeepSpeed-Megatron (Smith et al., 2022) propose custom 3D-parallelism strategies.
+DeepSpeed-ZeRO (Rajbhandari et al., 2020) and Pytorch-FSDP (Zhao et al., 2023) reduce redundant memory consumption in data parallelism.
+A set of works (Chen et al., 2024; Jiang et al., 2024; Chang et al., 2024; TE, 2024; Wang et al., 2022b) overlap the communication and computation operators in LLM training.
+These systems optimizations of LLM training are orthogonal to DistTrain.
+They do not address model and data heterogeneity in multimodal LLM training.
+DistTrain integrates several of these optimizations in training the LLM backbone.
+Traditional multimodal training.
+Traditional multimodal models (e.g., CLIP (Radford et al., 2021) and LiT (Zhai et al., 2022)) have been widely studied in recent years.
+Many systems optimizations have been proposed to train such multimodal models efficiently.
+DistMM (Huang et al., 2024) optimizes multiple parallel encoders training for contrastive learning.
+Yet, these advancements primarily enhance traditional multimodal models (e.g., CLIP (Radford et al., 2021) and LiT (Zhai et al., 2022)), which do not integrate LLM for understanding and generation (e.g., Flamingo (Alayrac et al., 2022) and LlaVa (Liu et al., 2024a)).
+They do not support encoder, LLM backbone and generator training pipeline in multimodal LLMs.
+Therefore, the state-of-the-art solution is to leverage Megatron-LM to train multimodal LLMs at large scale.
+10. Conclusion
+We present DistTrain, a disaggregated multimodal LLM training system.
+We identify the key challenges in training multimodal LLMs, i.e., model heterogeneity and data heterogeneity.
+DistTrain introduces disaggregated model orchestration to address model heterogeneity and disaggregated data preprocessing to address data heterogeneity.
+We evaluate DistTrain on a large production cluster with thousands of GPUs and show that it achieves 54.7% MFU and outperforms Megatron-LM by up to 2.2 on training throughput.
+This work does not raise any ethical issues.
+Acknowledgments.
+We sincerely thank the anonymous reviewers for their valuable feedback on this paper.
+Xin Jin is the corresponding author.
+Zili Zhang, Yinmin Zhong, and Xin Jin are also with the Key Laboratory of High Confidence Software Technologies (Peking University), Ministry of Education.
+This work was supported in part by the National Key Research and Development Program of China under Grant 2022YFB4500700, the Scientific Research Innovation Capability Support Project for Young Faculty under Grant ZYGXQNJSKYCXNLZCXM-I1, the Fundamental Research Funds for the Central Universities, Peking University, and the National Natural Science Foundation of China under Grant 62172008.
+Appendices are supporting material that has not been peer-reviewed.
+Appendix A Appendix
+A.1. Mitigating TP overhead with StepCCL
+Tensor Parallelism (TP) is commonly adopted to facilitate training large Transformer-based models with multiple GPUs connected with high bandwidth (e.g., NVLinks).
 Specifically, TP divides the linear layers, into smaller sub-modules, which are then distributed across the GPUs.
 After parallel computation, all GPUs perform collective communication to aggregate data and produce identical results.
-The TP communication overhead severely degrades overall performance, especially on A800 and H800 GPUs with restricted NVLink bandwidth.
+The TP communication overhead severely degrades overall performance.
 We implement the communication overlap with an in-house collective communication library called StepCCL to reduce the TP overhead.
 StepCCL is a PyTorch custom plugin that performs cross-GPU collective communication, including allgather, reduce-scatter, and allreduce, which is similar to NCCL.
-However, NCCL occupies several CUDA Streaming Multiprocessors (SMs) for executing its communication kernel and is known to harm the performance of its concurrent GEMM (matrix multiplication) [41].
+However, NCCL occupies several CUDA Streaming Multiprocessors (SMs) for executing its communication kernel and is known to harm the performance of its concurrent GEMM (matrix multiplication) (Rashidi et al., 2021).
 To solve this, StepCCL leverages the DMA engine directly to perform data transmission without using any SM at all.
@@
 Figure 21 describes the details of the layout remap process.
-In some rare cases during backward propagation, we find the remap overhead is high due to certain model dimension.
+In some rare cases during backward propagation, we find the remap overhead is high due to certain model dimensions.
 To mitigate this, we further overlap the remap with the computation of the weight gradients, so eventually we nearly get the full performance gain of the communication overlap.
-Finally, although the overlap idea is also studied in many related works [42, 43, 44], we highlight the key differences of ours.
-Unlike prior work that fuses GEMM with TP communication into one CUDA kernel [42, 43], we choose a modular design and do not use fusion for more flexibility.
+Finally, although the overlap idea is also studied in many related works (Chang et al., 2024; TE, 2024; Wang et al., 2022b), we highlight the key differences of ours.
+Unlike prior work that fuses GEMM with TP communication into one CUDA kernel (Chang et al., 2024; TE, 2024), we choose a modular design and do not use fusion for more flexibility.
 For example, when TP communication is longer than GEMM, fusing them cannot fully hide the communication overhead.
@@
 This enables broader adoption of StepCCL in many other scenarios.
-8 Evaluation
-In this section, we first use large-scale experiments to demonstrate the overall performance improvements of DistTrain over Megatron-LM.
-Next, we use microbenchmarks to deep dive into DistTrain and show the effectiveness of each component in DistTrain.
-Finally, we provide a case study to further elucidate DistTrain’s capabilities.
-| Models | # of | Hidden | FFN Hidden | # of | # of |
-| Layers | Size | Size | Heads | Groups | |
-| Llama3-7B | 32 | 4096 | 11008 | 32 | 32 |
-| Llama3-13B | 40 | 5120 | 13824 | 40 | 40 |
-| Llama3-70B | 80 | 8192 | 28672 | 64 | 8 |
-Setup.
-Our experiments are conducted on a production GPU cluster for multimodal LLM training, with each node equipped with eight NVIDIA A800 GPUs, 1TB of memory, and 96 vCPUs.
-GPUs within one node are interconnected by 300GB/s (bidirectional) NVLink, while nodes are connected by 4*200 Gbps RDMA network based on RoCEv2 with rail-optimized topology.
-The overall experiments use up to 1296 GPUs, and the microbenchmark utilizes up to 98 GPUs.
-We use PyTorch 2.1.2 and NVIDIA CUDA 12.2 for our evaluation.
-Models.
-For LLM backbone, we choose the representative LLM architecture, Llama3 [22], which is widely used in both academia and industry.
-Table 1 lists the detailed model configurations.
-As for the modality, we focus on images and texts.
-DistTrain is also compatible with other modalities.
-For modality encoder and modality generator, we use ViT-Huge [45] (0.63B) and Stable-Diffusion 2.1 [46] (1B) respectively.
-These two models are widely used for image understanding and generation.
-The three LLM backbones (i.e., Llama3-7B, Llama3-13B, and Llama3-70B) are paired with ViT-Huge and Stable-Diffusion to form multimodal LLMs designated as MLLM-9B, MLLM-15B, and MLLM-72B.
-For large multimodal LLM (i.e., MLLM-72B), we use high image resolution (i.e., 10241024) for generation since the large LLM is able to process more context information.
-For small models, we use low image resolution (i.e., 512512).
-Datasets.
-For our experiments, we use the representative open-source dataset, LAION-400M.
-We generate training data by interleaving the image and text subsequences, forming input sequences up to 8192 tokens long.
-This dataset is also employed in our production multimodal LLM training.
-As detailed in §2.3, each training sample includes a varying number of image tokens and text tokens, which introduces data heterogeneity in multimodal LLM training.
-Metrics.
-We use the Model FLOPs Utilization (MFU) as the primary metric to evaluate DistTrain.
-MFU measures the percentage of GPU FLOPs that are effectively utilized during model training.
-We also use the training throughput (TPT.) to evaluate the training speed of DistTrain.
-Since DistTrain and Megatron-LM may utilize different numbers of GPUs due to varying model orchestration strategies, we also indicate the number of GPUs used in the throughput charts.
-8.1 Overall Performance
-Setup.
-We first compare the overall performance of DistTrain against Megatron-LM on a large-scale GPU cluster (up to 1296 GPUs).
-We retrofit Megatron-LM to support multimodal LLM training by integrating modality encoder and generator into the training pipeline.
-Megatron-LM employs rigid model orchestration as described in §2.1.
-In Megatron-LM, we set the PP size of the LLM backbone to 1, 2, and 10 for Llama3-7B, Llama3-13B, and Llama3-70B.
-PP size is set to 1 for modality encoder and generator.
-TP size is set to 8.
-As for DistTrain, the parallelism strategy is determined by disaggregated model orchestration.
-In our experiments, one GPU is able to facilitate training ViT and Stable-Diffusion.
-We replicate the encoder and generator across the GPUs within the TP group to process different images, whereas TP itself is not used.
-We set global batch size to 1920.
-The experimental results are shown in Figure 15 and Figure 16.
-Figure 15 shows the MFU.
-Figure 16 shows the training throughput and marks the number of GPUs used in each experiment.
-The different GPU numbers are due to the varying model orchestration strategies of DistTrain and Megatron-LM.
-We summarize the experiment as follows.
-As shown in Figure 15, DistTrain achieves 51.8%-54.7% MFU in large-scale multimodal LLM training.
-This performance closely approximates that of state-of-the-art unimodal (i.e., text) LLM training [29], which demonstrates the effectiveness of DistTrain in addressing the model and data heterogeneity in multimodal LLM training.
-DistTrain significantly outperforms Megatron-LM, delivering 1.7-2.8 the MFU and 1.7-2.2 the training throughput when training MLLM-9B and MLLM-15B with a similar number of GPUs.
-These performance gains largely stem from DistTrain’s disaggregated model orchestration.
-Megatron-LM’s rigid strategy often leads to GPU underutilization, since it assigns too many GPUs to the modality encoder and generator.
-In contrast, DistTrain adaptively adjusts model orchestration based on specific model and data demands.
-Additionally, DistTrain’s disaggregated data reordering technique further boosts efficiency.
-In MLLM-72B training scenario, DistTrain also outperforms Megatron-LM by 1.2 on MFU and 1.3 on training throughput with a similar number of GPUs.
-The high image resolution prolongs the execution time of the multimodal module, which introduces pipeline bubbles in LLM backbone.
-DistTrain addresses this by allocating additional GPUs to these modules to balance the pipeline.
-The disaggregated data reordering strategy continues to diminish data heterogeneity, thereby increasing training efficiency.
-8.2 Deep Dive into DistTrain
-In this subsection, we perform microbenchmarks to evaluate the effectiveness of each DistTrain’s component.
-Our experiments utilize up to 98 NVIDIA A800 GPUs.
-We set the global batch size to 128, 64, and 40 for MLLM-9B, MLLM-15B, and MLLM-72B, respectively.
-8.2.1 Disaggregated Model Orchestration
-We measure the MFU and training throughput of DistTrain and other model orchestration strategies, including a comparison with Megatron-LM and a ratio-based approach.
-The ratio-based strategy allocates GPUs according to the computational demands (flops) of each module.
-We also annotate the total number of GPUs utilized in the throughput chart.
-The experimental results are shown in Figure 17.
-DistTrain consistently outperforms the baseline strategies, achieving 1.3-2.7 higher MFU and 1.4-2.7 higher training throughput.
-Although the ratio-based strategy outperforms Megatron-LM’s rigid strategy, it still lags behind DistTrain since it neglects the intricate performance model (§5.1) of multimodal LLM training.
-DistTrain’s disaggregated model orchestration optimally balances computational loads across the three modules and achieves high resource utilization.
-(a) MFU. (b) Throughput.
-We also evaluate the running time of DistTrain’s disaggregated model orchestration under different training settings, as detailed in Table 2.
-The algorithm completes in under one second.
-The overhead is negligible compared to the days or even weeks required for overall training.
-8.2.2 Disaggregated Data Reordering
-We evaluate the effectiveness of DistTrain’s disaggregated data reordering by comparing it against the random order, while keeping other components the same.
-The effectiveness is gauged through metrics such as MFU and training throughput (TPT.).
-We use the optimal resource allocation and parallelism strategy decided by DistTrain’s disaggregated model orchestration.
-Given that the model orchestration strategy remains unchanged, the number of GPUs is not shown.
-The experimental settings are the same as those in §8.2.1.
-The results are shown in Figure 18.
-DistTrain consistently outperforms the baseline, achieving 1.03-1.11 higher MFU and training throughput.
-The performance gap becomes more pronounced as the model size decreases.
-This is because the smaller model size leads to a higher data parallelism (DP) size, which causes more inter-microbatch heterogeneity.
-In essence, DistTrain’s disaggregated data reordering effectively mitigates data heterogeneity and enhances the training efficiency.
-We do not measure the running time of the reordering algorithm as it operates on dedicated CPU nodes asynchronously.
-It does not interfere with the GPU training process.
-The only overhead is the network delay which is evaluated in §8.3.
-| Model | # of | Global | Algorithm |
-| GPUs | Batch Size | Overhead | |
-| MLLM-72B | 1296 | 1920 | 922ms |
-| MLLM-72B | 648 | 960 | 641ms |
-| MLLM-72B | 324 | 480 | 441ms |
-| MLLM-72B | 112 | 240 | 133ms |
-8.3 Case Study
-In this subsection, we first evaluate DistTrain under different frozen training settings.
-We then evaluate the overhead of CPU preprocessing disaggregation and the effectiveness of StepCCL in mitigating the TP overhead.
-8.3.1 Frozen Training
-We conduct a frozen training experiment under four specific training settings: complete module freezing, exclusive encoder training, exclusive LLM training, and exclusive generator training.
-We keep training the projectors.
-In these scenarios, frozen modules neither compute weight gradients nor update weights.
-All other experimental setup aligns with those detailed in §8.2.
-We evaluate the MFU and training throughput (TPT.) of DistTrain compared to Megatron-LM.
-The experimental results are presented in Figure 19 and Figure 20.
-DistTrain consistently outperforms Megatron-LM across all frozen training configurations, achieving 1.4-2.9 higher MFU and 1.2-2.9 higher training throughput.
-This pronounced performance gap underscores the challenges posed by Megatron-LM’s rigid model orchestration in complex training environments.
-In contrast, DistTrain adaptively adjusts model orchestration based on training settings and consistently achieves high resource utilization.
-(a) MFU. (b) Throughput.
-(a) All modules freezing. (b) Exclusive encoder training. (c) Exclusive LLM training. (d) Exclusive generator training.
-(a) All modules freezing. (b) Exclusive encoder training. (c) Exclusive LLM training. (d) Exclusive generator training.
-Overhead of data preprocessing.
-We conduct an experiment to evaluate the overhead of data preprocessing, including decompression and reordering.
-Setting the DP size to one, we measure the average data preprocessing time per iteration on the GPU training side.
-We then compare data preprocessing time with and without CPU preprocessing disaggregation and use varying numbers of images and different image resolutions for one training iteration.
-The results, depicted in Figure 21, indicate that the disaggregation significantly reduces preprocessing time from seconds to milliseconds.
-The first parameter in the x-axis represents the number of images, while the second parameter denotes the image resolution.
-In production training (§8.1), iteration times range from seconds to tens of seconds.
-Preprocessing overhead, initially counted in seconds, significantly interferes with training.
-With disaggregated data preprocessing, the overhead reduces to milliseconds, which is negligible relative to total iteration time.
-Mitigating TP overhead with StepCCL.
+Evaluation.
 To evaluate the effectiveness of StepCCL in mitigating the TP overhead, we conduct an experiment that measures the iteration time of the LLM backbone with training of one single PP stage (i.e., one minimal TP group) under various TP sizes.
@@
 StepCCL significantly reduces the iteration time by overlapping the TP communication with computation.
-It outperforms the baseline by 1.1-1.17 when the TP size is 4 and 1.15-1.17 when the TP size is 8.
+It outperforms the baseline by 1.1-1.12 when the TP size is 4 and 1.15-1.17 when the TP size is 8.
 The gains are more pronounced at large TP size, where communication overhead is more substantial.
 These findings confirm that StepCCL effectively mitigates TP overhead.
-9 Discussion
-Parallelism optimization.
-Many studies [47, 48, 49] refine parallelism strategies for deep learning models, but they fall short for multimodal LLMs due to their expansive search spaces and large cluster requirements.
-These methods separately optimize different strategies, resulting in inefficient parallelism Additionally, these methods generally assume data homogeneity, standardizing computations across all training samples.
-This assumption does not hold for multimodal LLMs due to the data heterogeneity in modality inputs of each sample.
-In contrast, DistTrain leverages the specific training pattern of multimodal LLMs, creating a customized model orchestration problem that integrates tensor, pipeline, and data parallelism simultaneously.
-Besides, DistTrain leverages disaggregated data reordering to harness the data heterogeneity.
-Sequence and expert parallelism.
-Sequence parallelism (SP) [50] is designed to partition the training sequence into multiple subsequences for parallel training.
-It addresses the challenges of processing long sequences in LLMs.
-Expert parallelism (EP) [51], specifically devised for mixture-of-experts (MoE) LLMs [52], enables parallel training of multiple feedforward network (FFN) experts.
-These parallelism strategies are orthogonal to multimodal LLM training.
-In DistTrain, both SP and EP are integrated into the training framework.
-DistTrain treats SP and EP sizes as predefined parameters in the disaggregated model orchestration optimization problem.
 (a) TP=4. (b) TP=8.
-10 Related Work
-LLM training.
-Many efforts have been made to optimize the training of LLMs from system perspectives.
-For LLM pretrain, Megatron-LM [17] and DeepSpeed-Megatron [53] propose customized 3D-parallelism and are de facto standards for training large LLMs.
-DeepSpeed-ZeRO [32] and Pytorch-FSDP [54] reduce redundant memory consumption in data parallelism.
-HPN [55] proposes a new dual-ToR network architecture trailed for LLM training to reduce ECMP conflict.
-A set of works [56, 29, 42, 43, 44] overlap the communication and computation operators in LLM training.
-Fault tolerance through replication and checkpoint is advanced in large training clusters by studies [29, 57].
-Efforts like [58, 59, 60] further optimize recovery process in cloud spot instance scenarios.
-These system optimizations of LLM training are orthogonal to DistTrain.
-They overlook the model and data heterogeneity in multimodal LLMs.
-DistTrain also integrates several of these optimizations in training the LLM backbone.
-Multimodal model training.
-Small multimodal models (e.g., CLIP [61] and LiT [62]) have been widely studied in recent years.
-Many system optimizations have been proposed to train such multimodal models efficiently.
-DistMM [63] tackles model heterogeneity by introducing modality-aware placement and partitioning to evenly distribute workload.
-GraphPipe [64] presents graph pipeline parallelism, addressing graph dependencies in multimodal models to minimize pipeline bubbles.
-Yet, these advancements primarily enhance small multimodal models trained on tens of GPUs.
-They fall short for scaling up to meet the demands of integrating multimodal models with LLMs which necessitate training across thousands of GPUs.
-This gap underpins the motivation behind DistTrain, designed to meet the unique challenges of multimodal large language model training.
-Multimodal LLM serving.
-LLM serving has been widely studied in recent years.
-Orca [65], FastServe [66], and VTC [67] propose iteration-level scheduling algorithms to improve the serving quality.
-DistServe [68] and Splitwise [69] propose disaggregated serving for prefill and decoding to improve the serving throughput.
-vLLM [70], RAGCache [71], and SGLang [72] propose prefix caching to reuse the KV tensors.
-However, the serving of multimodal LLMs remains under-explored.
-DistTrain’s core insights (e.g., disaggregation) are applicable to multimodal LLMs serving.
-Our future work will delve into the specific challenges posed by multimodal LLMs serving systems.
-11 Conclusion
-We present DistTrain, an efficient and adaptive framework to reform the training of multimodal LLMs.
-We identify the key challenges in training multimodal LLMs, i.e., model heterogeneity and data heterogeneity.
-DistTrain introduces disaggregated training to address these challenges, including disaggregated model orchestration to address model heterogeneity and disaggregated data reordering to address data heterogeneity.
-We evaluate DistTrain on production cluster with thousands of GPUs and show that it achieves 54.7% MFU and outperforms Megatron-LM by up to 2.2 on throughput.